### v1

# Confidence Calibration in Vision-Language-Action Models

Abstract
Trustworthy robot behavior requires not only high levels of task success but also that the robot can reliably quantify how likely it is to succeed. To this end, we present the first systematic study of confidence calibration in vision-language-action (VLA) foundation models, which map visual observations and natural-language instructions to low-level robot motor commands. We begin with extensive benchmarking to understand the critical relationship between task success and calibration error across multiple datasets and VLA variants, finding that task performance and calibration are not in tension. Next, we introduce prompt ensembles for VLAs, a lightweight, Bayesian-inspired algorithm that averages confidence across paraphrased instructions and consistently improves calibration. We further analyze calibration over the task time horizon, showing that confidence is often most reliable after making some progress, suggesting natural points for risk-aware intervention. Finally, we reveal differential miscalibration across action dimensions and propose action-wise Platt scaling, a method to recalibrate each action dimension independently to produce better confidence estimates. Our aim in this study is to begin to develop the tools and conceptual understanding necessary to render VLAs both highly performant and highly trustworthy via reliable uncertainty quantification.
1 Introduction
Confidence calibration, or the degree to which a model’s predicted probabilities reflect the true likelihood of its predictions being correct, is a cornerstone of reliable machine learning systems (Guo et al., 2017). When a model is well-calibrated, downstream decision-makers (e.g., humans, planning algorithms, or safety monitors) can trust that a confidence estimate of 95% implies that the predicted outcome will occur roughly 95% of the time (see Figure 1). Mismatches between confidence and actual outcomes can have severe negative consequences in high-stakes settings such as medical diagnosis or autonomous driving, motivating a rich literature on improving calibration in deep learning models (Minderer et al., 2021; Lakshminarayanan et al., 2017; Gal and Ghahramani, 2016; Tian et al., 2023; Lin et al., 2024; Kadavath et al., 2022).
Figure 1: To be trustworthy, a robotic system must be able to reliably express its confidence in its ability to perform a task, especially in high-stakes and open-world domains. A well-calibrated robot policy produces confidence estimates that align with its probability of task success. For example, the robot should succeed on 95% of instances for which it expresses 95% confidence.
Recently, the field of robotics has embraced a new class of vision-language-action (VLA) foundation models (Zitkovich et al., 2023; Kim et al., 2025; Black et al., 2024, 2025). Leveraging large-scale, multistage pretraining, these models translate visual observations and natural language instructions into low-level joint-space commands, unifying multimodal perception with motor control. Although still a relatively new paradigm, current VLA systems already demonstrate previously unattainable generalization across environments, tasks, and robot embodiments. Because these models encode broad semantic and visuomotor priors from large-scale pretraining, they can be efficiently fine-tuned to new robots and downstream tasks, yielding large gains over training from scratch.
Since VLAs are designed for closed-loop interaction with the physical world, knowing when and how strongly to trust their actions is critical. For example, consider a robot performing a task in a safety-critical environment, or attempting to manipulate a valuable, fragile object. In these scenarios, if the policy can accurately express uncertainty, then costly or dangerous accidents can be avoided, such as by refining the instruction or deferring the task to a human. Despite the importance of calibrated confidence, basic questions remain largely unaddressed, e.g., whether VLAs are calibrated, how calibration evolves over the task horizon, and how it can be improved.
Contributions. To bridge this gap, we present (to our knowledge) the first systematic study of calibration in VLAs, identifying key open questions and introducing practical remedies for miscalibration. Our contributions include:
  1. 1.
    Benchmarking task success vs. calibration. We perform a comprehensive evaluation to measure the relationship between task success and calibration error across multiple datasets and VLA variants, finding that better task performance is associated with better calibration.
  2. 2.
    Prompt ensembles. We propose a lightweight, Bayesian-style method that averages a VLA’s confidence across multiple semantically equivalent rephrasings of an instruction. This approach consistently improves calibration, cutting expected calibration error by more than 20% on average.
  3. 3.
    Calibration over task time. We analyze calibration over task time, showing that confidence is often most reliable after making some task progress, suggesting natural points for risk-aware intervention.
  4. 4.
    Action-wise scaling. We discover systematic over-/underconfidence in different action dimensions and propose a method to recalibrate each action dimension independently to produce more reliable confidence estimates.
Our aim in this study is to begin to develop the tools and conceptual understanding necessary to render VLAs not only highly performant but also highly trustworthy via reliable uncertainty quantification.
2 Related Work
2.1 Confidence Calibration
A model is considered well-calibrated when the confidence (i.e., probability) it assigns to an outcome matches the long-run frequency of that outcome. Deviation from this condition is often quantified via expected calibration error (ECE) (Guo et al., 2017), which is approximated by grouping predictions with similar confidence into bins and measuring the absolute difference between confidence and accuracy in each bin. Beyond ECE, other metrics such as maximum calibration error (Guo et al., 2017), Brier score (Brier, 1950), and negative log-likelihood (NLL) are used to capture related and complementary notions of miscalibration. Despite their impressive accuracy and trends toward better calibration, modern neural networks still display some persistent miscalibration (Minderer et al., 2021), especially when the task distribution does not match the training distribution (Ovadia et al., 2019). Consequently, a rich toolbox of post hoc fixes has emerged: Platt scaling (Platt, 1999), temperature scaling (Guo et al., 2017), histogram binning (Zadrozny and Elkan, 2001), and other recalibration methods (Kumar et al., 2019; Naeini et al., 2015; Zollo et al., 2024) can be applied to an already trained model to shrink residual calibration error without altering its decision rule. Other popular methods to improve calibration of neural networks include dropout (Gal and Ghahramani, 2016) and ensembling (Lakshminarayanan et al., 2017).
2.2 Vision-Language-Action Models
Vision-Language-Action models take visual data and natural language instructions as input and output robot actions (Zitkovich et al., 2023; Kim et al., 2025; Black et al., 2024, 2025). They are typically initialized from a visually-conditioned language model (VLM) that, in turn, is initialized from a pretrained large language model (LLM). This allows VLAs to leverage rich multimodal priors in joining perception and action generation into a single end-to-end pipeline. While some VLAs retain the token-based output paradigm inherited from these base models (Zitkovich et al., 2023; Kim et al., 2025), others have augmented the architecture with, e.g., flow matching action experts for smooth, high-frequency control (Black et al., 2024, 2025). These systems already perform complex, language-specified tasks across diverse environments and robot embodiments. However, they lack a reliable mechanism for quantifying the uncertainty of their chosen action sequences.
2.3 Calibration in LLMs
Given that most state-of-the-art VLAs are built on LLM backbones, it is natural to consider how our work relates to existing studies of calibration in LLMs, an area that has received considerable attention recently. Typical approaches to generating confidence estimates with LLMs include expressing calibration as a multiple-choice question for the LLM (Kadavath et al., 2022), directly verbalizing confidence with the model’s text output (Lin et al., 2022; Tian et al., 2023; Band et al., 2024), or measuring semantic consistency across many sampled outputs for the same query (Kuhn et al., 2023; Duan et al., 2024; Chen et al., 2024). However, it is difficult to adapt these methods to the VLA setting. For example, while LLMs may produce the entire sequence before measuring confidence, in robotics potential failures must be flagged much earlier in the trajectory to ensure safety and avoid costly accidents. Also, sampling multiple full trajectories in the physical world may be impossible. Finally, current VLAs lack the flexible and robust text-to-text interface of LLMs, and thus cannot be expected to, e.g., answer natural language questions about their confidence in an action prediction.
2.4 Uncertainty Quantification in Robotics
Uncertainty quantification has long been a focus in robotics, particularly through the lens of probabilistic robotics (Thrun et al., 2005; Deisenroth and Rasmussen, 2011). Recent work explores the use of model ensembles in control systems, examining how disagreements among multiple models can capture uncertainty and guide more cautious policy updates and actions (Chua et al., 2018; Pathak et al., 2019; Ataei and Dhiman, 2024). Other advances include uncertainty-based active exploration (Wang et al., 2024) and the application of conformal methods, which can provide distribution-free guarantees for planning or classification pipelines (Ren et al., 2023; Sun et al., 2023). How to obtain calibrated, task-level uncertainty estimates given the expressive, multimodal perception and control abilities of VLAs remains an open and safety-critical problem.
3 Calibration in Vision-Language-Action Models
In this section, we formalize the problem of confidence calibration for vision-language-action (VLA) policies. We define calibration, describe how contemporary VLA architectures generate actions and how to extract a corresponding confidence estimate, and introduce the standard calibration metrics used in our empirical study. Finally, we present two lightweight remedies for the miscalibration we observe in practice: prompt ensembles and action-wise Platt scaling.
3.1 Calibration
Let $C\!\in[0,1]$ denote the confidence reported by a robot policy and $Y\!\in\{0,1\}$ the binary indicator of task success (we use uppercase $C,Y$ for random variables and lowercase $c,y$ for their realizations). A perfectly calibrated predictor satisfies
$\mathbb{P}(Y=1\mid C=c)\;=\;c,\qquad\forall\,c\in[0,1].$ (1)
If this condition is met, then for the subset of trials on which the robot reports 80% confidence, we should observe successful completion 80% of the time.
Figure 2: Given an input image and text instruction, popular VLAs such as OpenVLA and RT-2 generate a distribution over discrete action tokens for each of the robot’s degrees of freedom. Confidence in each dimension’s prediction can be estimated using the probability assigned to the predicted token; a single estimate can be produced by averaging across dimensions. Given an uncalibrated confidence estimate, recalibration methods such as Platt scaling use a small calibration dataset to learn a map from uncalibrated confidence estimates to calibrated ones. Calibration error can be measured by comparing confidence estimates to actual task success rates.
3.2 Vision-Language-Action Models
At task timestep $t$, a VLA policy $\pi_{\theta}$ has access to the observed history $o_{t}=(v_{\leq t},l_{\leq t},a_{<t})$, where $v$ is a visual observation, $l$ is the (possibly fixed) natural language instruction, and $a$ is an action. The policy induces a distribution over the next action $\pi_{\theta}(a\mid o_{t}).$
When performing a task, the policy executes the most likely action $a_{t}^{\star}=\mathop{\rm argmax}_{a}\pi_{\theta}(a\mid o_{t})$. Given the need for calibration, we extract a scalar confidence score $c_{t}$ from the policy. We interpret $c_{t}$ as the policy’s estimate of the probability that the task will ultimately succeed given the observation history and the chosen action, i.e., $\mathbb{P}(Y=1\mid o_{t},a_{t}^{\star})$. Our subsequent analysis and metrics treat $c_{t}$ as a black-box number, so the following sections apply equally to token-based, diffusion-based, or other controllers, as long as they are able to emit such a scalar.
3.3 Baseline Confidence Estimation
Many state-of-the-art VLAs, including OpenVLA (Kim et al., 2025) and RT-2 (Zitkovich et al., 2023), represent an action with $D$ discrete tokens, where each represents one dimension of the robot’s action space. For each dimension $d\in\{1,\dots,D\}$ the policy outputs logits $z^{(d)}_{t}$ and probabilities
$p^{(d)}_{t}=\operatorname{softmax}\!\bigl{(}z^{(d)}_{t}\bigr{)},\qquad p^{(d)}_{t,k}=\pi_{\theta}\!\bigl{(}A^{(d)}_{t}=k\mid o_{t}\bigr{)}$ (2)
for action tokens $k\in\{1,\dots,K\}$ in an action vocabulary of size $K$. At time $t$, the policy selects the top tokens $a^{(d)}_{t}=\mathop{\rm argmax}_{k}p^{(d)}_{t,k}$ for each action dimension $d$, and decodes them into a continuous action for execution by the robot.
Because VLAs with token-based decoders closely mirror LLMs, we can adopt the usual LLM heuristic of using the probability of the token actually chosen as a confidence signal. For each action dimension, we take the probability assigned to the selected token, and average these values:
$c_{t}\;=\;\frac{1}{D}\sum_{d=1}^{D}\max_{k}p^{(d)}_{t,k}.$ (3)
Averaging plays the same role as length-normalization in LLMs, preventing confidence from being unfairly reduced for robots with many degrees of freedom. We use this as our baseline confidence estimate in our experiments.
3.4 Measuring Calibration
Calibration metrics translate deviations from the condition in Equation (1) into quantitative measures of miscalibration. Given the difficulty of measuring such a condition (i.e., comparing two distributions), we consider a range of metrics.
For what follows, let $\{(c_{i},y_{i})\}_{i=1}^{N}$ denote the reported confidence and binary outcome for each of $N$ robot trials (episodes). Each trial $i$ consists of timesteps $t=1,\ldots,T_{i}$ with per-timestep confidences $c_{i,t}$; to obtain a single trial-level value $c_{i}$ we apply an aggregation function $h$: $c_{i}=h(\{c_{i,t}\}_{t=1}^{T_{i}})$. Possible choices include the confidence before the first action, the mean across timesteps, or the min/max over the trajectory. The following measures are agnostic to choice of $h$; in our experiments we use the pre-action confidence $c_{i}=c_{i,1}$, reflecting the high-stakes open-world robotics setting in which early risk assessment is particularly valuable.
Expected Calibration Error.
One popular measure of miscalibration is expected calibration error (ECE) (Guo et al., 2017; Fisch et al., 2022):
$\text{ECE}_{q}\;=\;\bigl{(}\mathbb{E}_{C}\!\bigl{[}\lvert\mathbb{P}(Y=1\mid C)-C\rvert^{q}\bigr{]}\bigr{)}^{\!1/q}\!.$ (4)
Put simply, ECE measures the expected difference between confidence and accuracy over the robot’s task data distribution. The parameter $q$ is typically set to $q\in\{1,2\}$, where $ECE_{1}$ weights deviations linearly and $ECE_{2}$ penalizes larger errors.
Because we can observe only a finite sample of trial results, the conditional expectation in Equation (4) cannot be directly measured. Instead, it is typically approximated with a binning-based estimator. With results from $N$ robot trials, we approximate the population ECE quantity in Equation (4) by first ordering predictions according to confidence, and splitting them into $M$ equal-sized bins $B_{1},\dots,B_{M}$. Then, our empirical estimate $\widehat{\text{ECE}}_{q}$ is given by
$\widehat{\text{ECE}}_{q}\;=\;\Bigl{(}\sum_{m=1}^{M}\frac{|B_{m}|}{N}\bigl{|}\mathrm{acc}(B_{m})-\mathrm{conf}(B_{m})\bigr{|}^{q}\Bigr{)}^{\!1/q},$ (5)
where
$\mathrm{acc}(B_{m})=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}y_{i},~{}~{}~{}~{}\mathrm{conf}(B_{m})=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}c_{i}.$
We report both $\text{ECE}_{1}$ and $\text{ECE}_{2}$ in our experiments.
Brier Score.
Brier score (Brier, 1950) is a classic measure of the quality of probabilistic forecasting:
$\mathrm{BS}=\frac{1}{N}\sum_{i=1}^{N}(c_{i}-y_{i})^{2}.$ (6)
Brier score is an example of a proper scoring rule (Gneiting and Raftery, 2007), meaning that it is minimized only when the predicted distribution matches the true distribution for each example (as opposed to ECE, where data is grouped by confidence level). Brier score rewards both reliability (confidence matching success) and sharpness (predictions away from the population rate).
Negative Log-Likelihood.
Another metric used to measure calibration (Guo et al., 2017) is negative log-likelihood (NLL):
$\mathrm{NLL}=-\frac{1}{N}\sum_{i=1}^{N}\!\bigl{[}y_{i}\log c_{i}+(1-y_{i})\log(1-c_{i})\bigr{]},$ (7)
Also a proper scoring rule, NLL penalizes very confident failures much more heavily than Brier score.
Complementary Roles of Metrics.
Because each metric emphasizes a different facet of probabilistic quality, we report all three to obtain a detailed view of calibration performance. ECE focuses on reliability, or the alignment between stated confidence and empirical success, but is indifferent to how decisive those confidences are. A trivial predictor that outputs the base success rate on every trial achieves zero ECE while offering almost no utility for decision-making; Brier score and NLL would penalize this behavior. Conversely, since $Y$ is binary and only observed once, both Brier score and NLL implicitly discourage probabilistic predictions in the interior of $[0,1]$ when the goal is to achieve a very low loss. However, such estimates may sometimes be appropriate, e.g., due to randomness in the robot’s environment. Evaluating all three metrics together thus guards against corner cases and ensures that improvements in controlled evaluations translate into more actionable and reliable confidence estimates that are simultaneously calibrated, discriminative, and informative.
3.5 Prompt Ensembles
Similar to the VLMs from which they are derived (Zhou et al., 2024), semantically meaningless lexical differences in instructions, e.g., “pick up the coffee cup” vs. “grab the mug”, might shift visual attention, alter path planning, and thereby change the predictions and confidence scores emitted by VLA models. Although such sensitivity can pose a challenge to reliable task execution, it also creates an opportunity to employ an ensemble-based approach (Lakshminarayanan et al., 2017) to confidence estimation. Specifically, we can treat the particular wording of an instruction as a latent random variable and marginalize over it via Bayesian model averaging. Such averaging over rephrasings reduces variance in the final confidence estimate by canceling out the noise induced by word choice. Implementing this idea as an algorithm, we can then:
  1. 1.
    Generate rephrasings. An auxiliary LLM produces $r$ semantically equivalent prompts $\mathcal{L}_{alt}=\{l_{alt}^{(1)},\dots,l_{alt}^{(r)}\}$ (see Table 1 for example).
  2. 2.
    Estimate confidence with each variant. Generating an action with $l_{alt}^{(i)}$ yields a confidence $c_{t}^{(i)}$ via Equation (3).
  3. 3.
    Aggregate. The final estimate is the ensemble mean $c^{ens}_{t}=\frac{1}{r}\sum_{i=1}^{r}c_{t}^{(i)}$.
Conceptually, this prompt ensemble technique can play a similar role to full model ensembles (Lakshminarayanan et al., 2017) or inference-time dropout (Gal and Ghahramani, 2016). With batched inference, this method should incur negligible latency with added inference time complexity $\mathcal{O}(1)$.
Instruction Pick up the black bowl between the plate and the ramekin and place it on the plate
Rephrasings (1) Lift the black bowl located between the plate and the ramekin and set it on the plate.
(2) Grasp the black bowl found between the plate and the ramekin and move it to the plate.
(3) Take the black bowl positioned between the plate and the ramekin and position it on the plate.
Table 1: Example of multiple paraphrases of a single VLA instruction (from our experiments with LIBERO Spatial). The rephrasings carry the same meaning as the original instruction, enabling a reprompting approach in which we ensemble over predictions conditioned on lexically different yet semantically equivalent instructions.
3.6 Action-Wise Scaling
A standard remedy for miscalibration in classification models is to gather a small validation set from the task distribution and perform post hoc recalibration, learning a function that maps uncalibrated confidence estimates to calibrated ones (Platt, 1999; Naeini et al., 2015; Zadrozny and Elkan, 2001; Guo et al., 2017). Post hoc recalibrators are typically simple functions of the original confidence estimate.
One popular example of a post hoc recalibration method is Platt scaling (Platt, 1999; Kumar et al., 2019). Given confidence outputs and binary outcomes $\{(c_{i},y_{i})\}_{i=1}^{N}$ on a held-out validation set, one fits an affine transform $g(c)=\sigma(\alpha c+\beta)$ that minimizes NLL,
$\min_{\alpha,\beta}-\sum_{i}\Bigl{[}y_{i}\log g(c_{i})+(1-y_{i})\log\!\bigl{(}1-g(c_{i})\bigr{)}\Bigr{]},$
where $\sigma(x)=1/(1+e^{-x})$. At inference each $c_{t}$ is replaced with $\tilde{c}_{t}=g(c_{t})$, with the goal of aligning confidence with the probability of task success while leaving the model’s actions unchanged.
Unlike classification models that emit one predictive distribution per sample, token-based VLAs output one predictive distribution per action dimension. Those dimensions can differ dramatically, for example because gripper open/close appears in nearly 100% of demonstrations, whereas “rotate wrist” is rarer. A single global transform therefore may not be able to correct all dimensions simultaneously.
Action-Wise Platt Scaling. We propose to address this heterogeneity by fitting one affine transform per dimension $d$:
$\tilde{c}_{t}\;=\;\frac{1}{D}\sum_{d=1}^{D}\sigma\!\bigl{(}\alpha_{d}\,\max_{k}p^{(d)}_{t,k}+\beta_{d}\bigr{)},$ (8)
with parameters $\{\alpha_{d},\beta_{d}\}$ learned on the same calibration set. Intuitively, $\alpha_{d}$ scales dimension $d$, flattening or sharpening its distribution, while $\beta_{d}$ shifts its overall optimism. As in standard Platt scaling, the chosen tokens $a^{(d)}_{t}$ remain unchanged; we only modify the reported confidence.
This dimension-wise perspective can be applied to other post hoc recalibrators as well, for instance action-wise temperature scaling. Beyond the specific methodology, we aim to highlight that VLA calibration will require domain-specific tools rather than a direct transplant of methods designed for standard classifiers.
4 Experiments
Having described the problem of confidence calibration in vision-language-action models, we next conduct an extensive empirical investigation, studying the following questions:
  1. 1.
    How are task success and calibration error related? Does a higher success rate imply better calibration? (Section 4.2)
  2. 2.
    Can prompt ensembles consistently improve confidence estimates? (Section 4.3)
  3. 3.
    How does calibration evolve over the task time horizon? (Section 4.4)
  4. 4.
    Are action dimensions differentially calibrated, and what are the implications for recalibration? (Section 4.5)
Our aim is to highlight key issues and lay the empirical groundwork for future research on calibrating VLAs. [^1] Our code is available at https://github.com/thomaspzollo/vla_calibration.
4.1 Experiment Details
We perform our experiments using models derived from the OpenVLA model (Kim et al., 2025). In particular, we use models fine-tuned on 3 different task suites from the LIBERO (Liu et al., 2023) benchmark: Spatial, Object, and Goal. LIBERO is a simulation environment for language-conditioned robot manipulation tasks inspired by human activities (see Table 2 for task examples). We use 3 different fine-tuned models for each task suite: fine-tuned OpenVLA, as well as its 8-bit and 4-bit quantized versions.
Each task suite features 10 different tasks with 50 randomized initializations, for a total of 500 examples. For a given task suite, task success rate represents the proportion of the 500 trials that result in success (while task error rate represents the proportion of trials that result in failure). In order to contextualize calibration results, task success rates are reported in Table 3. To calculate ECE, we use 12 equal-mass bins and the Python package released with Kumar et al. (2019).
Unless otherwise noted, we focus on confidence estimates produced at the first timestep, before any action is taken. This aligns with the need for safety in open-world robot deployments, where robots should signal uncertainty as early as possible in order to avoid costly or dangerous incidents. In Section 4.4, we explore how calibration differs across timesteps over the task horizon.
Dataset Example Image Example Instruction
Spatial [Uncaptioned image] Pick up the black bowl between the plate and the ramekin and place it on the plate.
Object [Uncaptioned image] Pick up the alphabet soup and place it in the basket.
Goal [Uncaptioned image] Open the middle drawer of the cabinet.
Table 2: Task examples from the LIBERO Spatial, Object, and Goal task suites.
Model Spatial Object Goal
OpenVLA (Full) 0.828 0.880 0.758
OpenVLA (Quant-8) 0.844 0.876 0.760
OpenVLA (Quant-4) 0.822 0.850 0.712
Table 3: Task success rates across all 3 fine-tuned versions of OpenVLA and 3 different LIBERO task suites.
Figure 3: Visualization of task error rates compared against 4 different calibration error measurements for all model variants and task suites. Lower task error rates are associated with lower calibration error.
4.2 Relationship Between Calibration and Task Success
Much of the early research on calibration in deep learning models was based on perceptions of how task success (usually image classification accuracy) was related to calibration error. In particular, influential work argued that modern neural networks are poorly calibrated (Guo et al., 2017), suggesting that improved accuracy comes at the expense of calibration. This led researchers to propose training interventions to augment the cross-entropy loss with other objectives seeking improved calibration, potentially at a cost to overall accuracy (Kumar et al., 2018; Mukhoti et al., 2020). However, subsequent research found that, in fact, more accurate networks are generally better calibrated (and easier to recalibrate), and thus modifications to training procedures may not be needed (Minderer et al., 2021). Instead, techniques such as post hoc recalibration (Guo et al., 2017; Zadrozny and Elkan, 2001; Kumar et al., 2019) and ensembling (Lakshminarayanan et al., 2017; Fort and Lakshminarayanan, 2024), applied to models trained for high accuracy, are sufficient to achieve low calibration error (although calibration under distribution shift remains a significant challenge (Ovadia et al., 2019)).
To establish high-level direction for calibration research in VLAs, our first experiment focuses on this important question of how task success relates to calibration error. To study this question, in Figure 3 we plot calibration error (according to 4 different metrics) against task error rate for each of 9 combinations of 3 task suites and 3 model versions. Confidence estimates are produced using the baseline method described in Section 3.3 (a similar result under the prompt ensemble technique is shown in Appendix Figure 10).
Across all plots, a consistent trend emerges: lower task error rates are associated with lower calibration error. This mirrors findings in other areas of deep learning (e.g., as shown in image classification by Minderer et al. (2021)), and suggests that advances in task performance can naturally bring improvements in calibration. In that case, research can be focused on developing appropriate post hoc recalibration and ensembling techniques, as well as other bespoke solutions that can complement the simple and scalable learning paradigms that dominate modern deep learning.
4.3 Ensembling Confidence Over Prompts
We next study the empirical effectiveness of the prompt ensemble approach described in Section 3.5. First, for the natural language instruction associated with each task, we create 20 rephrasings using GPT-4o-mini (see Appendix Table 5 for prompts used). During testing, we produce a confidence estimate conditioned on each “reprompt”, and average over these confidence scores to obtain the final model confidence. Results are reported across 3 model variants and 3 task suites, and we measure $\text{ECE}_{1}$, $\text{ECE}_{2}$, Brier score, and NLL, comparing to the baseline method for producing confidence estimates using the original instruction.
Detailed results are recorded in Table 4 (results for the 4-bit quantized model are deferred to Appendix Table 7), while Figure 4 offers a visual summary of the high-level trends. Across all models, datasets, and metrics (36 measurements in total), the Reprompt (prompt ensemble) approach always decreases calibration error. Decreases in ECE can reach up to 40%, and average decreases are more than 20% (see Figure 11 for a visualization of percent changes across all metrics). The improvement is larger for ECE than the proper scoring rules (Brier score and NLL), where average miscalibration decreases are closer to 5-10%. This suggests that improvements in reliability (i.e., confidence matching marginal success rates) are generally greater than those in sharpness, which is expected from an ensemble technique targeted at reducing variance in confidence estimation. Overall, given its lightweight nature and strong empirical effectiveness, such data augmentation approaches seem like a promising direction for enhancing uncertainty quantification in VLAs.
Model Dataset Method $\text{ECE}_{1}$ $\text{ECE}_{2}$ Brier score NLL
OpenVLA Spatial Baseline 0.088 0.106 0.150 0.533
Reprompt 0.052 0.068 0.145 0.477
Object Baseline 0.060 0.073 0.108 0.401
Reprompt 0.039 0.053 0.105 0.362
Goal Baseline 0.151 0.170 0.207 0.707
Reprompt 0.115 0.132 0.197 0.620
Quant-8 Spatial Baseline 0.070 0.093 0.138 0.488
Reprompt 0.050 0.062 0.131 0.434
Object Baseline 0.057 0.065 0.110 0.409
Reprompt 0.041 0.052 0.108 0.375
Goal Baseline 0.140 0.163 0.205 0.713
Reprompt 0.117 0.152 0.194 0.608
Table 4: Calibration error measurements for the base and 8-bit fine-tuned model versions and 3 different LIBERO task suites, for 2 different methods of confidence estimation: (1) baseline (average selected token probability) (2) Reprompt (ensembling over semantically equivalent prompts). The prompt ensemble method improves all measures across all settings.
Figure 4: Comparing calibration error measurements across 2 different methods of confidence estimation: (1) baseline (average selected token probability) (2) Reprompt (ensembling over semantically equivalent prompts).
Figure 5: Ablation results for the prompt ensemble method, where different prompts are given to GPT-4o-mini for producing the rephrasings of the original instruction. Prompt 1 is used in the main experiments, while Prompts 2 and 3 are variants. Results show that improvements in calibration error from Reprompt are robust to different rephrasing prompts.
Figure 6: Ablating the number of prompts in the prompt ensemble for the Spatial task suite. Increasing the number of prompts generally improves calibration error.
4.3.1 Robustness Ablations for Prompt Ensembles
To understand the robustness of these results, we perform multiple ablations. First, we consider the effect of changing the prompt given to GPT-4o-mini for producing the 20 instruction rephrasings. In addition to the prompt used to produce the results in Table 4, we try two additional prompts, recorded in Appendix Tables 5 and 6; “Prompt 1” is the original prompt, and “Prompt 2” and “Prompt 3” represent new prompts for the ablation. Note that the different prompts lead to substantially different rephrasings, as Prompts 2 and 3 mandate the retention of the words “pick” and “place”, while as Table 1 shows, Prompt 1 leads to these words being replaced. The ablation is run across all 3 task suites using fine-tuned OpenVLA, and we measure $\text{ECE}_{1}$ and NLL. Results are shown in Figure 5. Improvements in calibration error from the Reprompt method are robust to different rephrasing prompts, with all ensemble variants producing lower error than the baseline method across both metrics and all 3 task suites.
Second, we consider the effect of the number of prompts used in the prompt ensemble. Given a sound ensembling approach, we would expect to see improvement as more prompts are added to the ensemble (up to some point). To study this question, we record results using $k\in\{1,5,10,20\}$ different rephrasings (randomly chosen for 1000 trials when $k<20$). Results for the Spatial task suite are shown in Figure 6, and additional results for the Goal and Object task suites are in Appendix Figure 12. We see that the algorithm behaves favorably, where calibration error generally decreases as more instructions are included in the ensemble.
4.4 Calibration Over Task Time
The previous experiments evaluate confidence before the first action is executed, a conservative choice for safety-critical deployments. Yet many tasks might allow the robot to collect more information without risk before having to express confidence. For instance, in the wine glass scenario of Figure 1, the gripper can hover above the stem, refine its scene representation, and only then decide whether it is confident enough to proceed. More context should, in principle, yield better calibration.
To test this intuition we measure calibration across 500 test trials for 100 different levels of task completion ($\{0,1,\dots,99\}\%$), where task completion is calculated as the current timestep index $t$ divided by the total number of timesteps in the task episode (and multiplied by 100). Our goal is to understand whether the quality of confidence estimates improves as the robot progresses in its task. At each measured timestep, we compute $\text{ECE}_{1}$ and Brier score using the baseline method for producing confidence estimates. Because a downstream safety monitor might consider the history of recent estimates (beyond the current one), we evaluate three aggregation rules for reporting a confidence estimate at each timestep:
  1. 1.
    Current - confidence from the current timestep only;
  2. 2.
    Window (5) - mean over the baseline confidence estimates from the current timestep and the four immediately preceding ones;
  3. 3.
    Avg. All - mean over the baseline confidence estimates from all steps seen so far.
To better understand the calibration error results, we also visualize how average confidence estimates evolve over the task horizon, separated by successful and failed trials. Finally, we plot reliability diagrams for the timesteps corresponding to $\{0,50,75,99\}\%$ completion. Reliability diagrams (Guo et al., 2017) offer a visualization of expected calibration error, using a similar binning strategy. They show confidence against accuracy for each bin, where a well-calibrated system lies on the $x=y$ line. Results for the Spatial task suite with the fine-tuned OpenVLA model are shown in Figure 7. Additional results for Spatial with the Quant-8 model (Figure 13), as well as the Object and Goal suites with the full precision (Figures 14 and 16) and Quant-8 models (Figures 15 and 17), are presented in Appendix B.3.
Figure 7: Empirical study of calibration error across the task time horizon for the Spatial task suite and the full precision fine-tuned OpenVLA model. In the top row, the left two plots characterize how calibration error evolves with task progress, while the right two plots show the average confidence estimates by task time, grouped by successful and failed trials. The bottom two rows offer a sample of the reliability diagrams produced by different methods for aggregating confidence estimates across time. Overall, these results illustrate that calibration can improve as the task progresses and more information is gathered, suggesting opportunities for context-aware uncertainty interventions.
Focusing first on results using the confidence estimate from the current step, a clear pattern emerges. Across all 6 settings (3 task suites each with 2 model variants) and both metrics, calibration improves sharply from 0% to approximately 50% completion, then plateaus or deteriorates back towards the original level (left-most and left-middle plots in the top row of the results figures). Beyond lower scores on calibration error, our additional plots characterize the improved probabilistic nature of the confidence estimates towards the middle of the task horizon. In the right-middle and right-most plots in the top row (titled “Successful Trials” and “Failed Trials”), the gap between the average confidence on successful vs. failed trials tends to be greatest around 50% task completion. Additionally, considering the top row of reliability diagrams, we can see that the confidence estimates using the current step become far more reliable around this point: error for many bins is large at the beginning, then the difference between confidence and accuracy for most bins becomes smaller until the task is roughly 75% complete. Near task completion, the gap between confidence on successful vs. failed trials closes again, and reliability suffers as a result.
These results support our hypothesis regarding the utility of gathering more information before expressing a confidence estimate, even using the simple baseline method. Early miscalibration might be expected: with only a high-level image frame and a textual instruction, the policy lacks much information about how hard the task might be. By mid-trajectory, it might observe how its gripper and the scene interact, analogous to a “test grip” on the wine glass, so its probability estimates may better match reality. Near the goal state, however, calibration error is generally worse, for instance because of differences in the action dimensions being emphasized in each phase of the task. These experiments suggest a practical recipe: let the robot execute a certified-safe prefix of the trajectory, and then assess its confidence and intervene if necessary. Such horizon-aware monitoring balances safety (no contact made during the prefix) with the improved reliability in confidence estimates that comes from extra perceptual context.
Beyond the baseline strategy of estimating confidence based only on the current timestep, we next observe the effects of other simple strategies for aggregating across timesteps. Specifically, we consider a sliding window average over 5 single-step estimates, and the average of all observed single-step estimates. Across all model/task suite combinations, the sliding window method is particularly effective at improving the Brier score around the middle of the task horizon, although ECE is not always improved. With respect to averaging all single-step estimates so far, we observe that Brier score improves throughout the task, but ECE generally gets worse. The underlying behavior driving these changes can be observed in the bottom row of reliability diagrams: the “Avg. All” method is fairly successful at sorting the failed examples to have relatively lower confidence, but also reduces variance in the confidence estimates such that they are highly overconfident in this range. Thus, while this approach fares poorly according to ECE, it is actually promising in the sense that it enables more effective discrimination between successful and failed trials.
4.4.1 Qualitative Examples of Context-Aware Confidence Monitoring
To illustrate how context-aware confidence monitoring could work in practice, we apply it to a representative pick and place task from the Goal suite. The task is to “put the wine bottle on the rack”, a case where the robot should be relatively conservative to avoid breaking glass. We consider a naive approach to context-aware monitoring, proposing to halt task performance when both:
  1. 1.
    The confidence level falls below a threshold set to the 10% quantile of confidence estimates for that point in the task horizon across all task trials (based on percent completion).
  2. 2.
    The robot is within a few inches of contacting an object, or already has contacted an object.
Since building a system to detect proximity to objects is beyond the scope of this work, we focus on a qualitative demonstration of this idea. For each example under examination, we plot current confidence for $\{0,20,40,60,80,98\}\%$ completion, as well as the corresponding 10% quantile risk threshold and an image of the robot environment at that time.
Some particularly illustrative examples are shown in Figure 8. First, we consider the episode shown across the top row (Episode 29 of the 50 in the Goal task suite). Here, confidence begins high, but falls until it is slightly below the halting threshold at 20% task completion, coinciding with the difficult subtask of gripping the rounded glass bottle. Allowed to continue, the model is able to grip the bottle, and confidence rebounds as it approaches the wine rack. However, confidence once again falls quickly when it begins to set the bottle down, possibly because the difficult grip put the bottle in an unfavorable position for placement. In the end, the robot fails to place the bottle securely on the rack, with potentially negative real-world consequences. Given the opportunity, a safety intervention could have been performed at multiple points before this incident, either to avoid grasping the bottle in the first place or to reset the bottle safely on the table instead of trying a failed placement.
Figure 8: Qualitative examples of a context-aware confidence monitoring strategy applied to a task from the Goal suite. Here, the task is to “put the wine bottle on the rack”. The red dashed line represents the 10% quantile of the confidence estimates output by the model across the task time horizon, offering a potential threshold below which the robot may abstain from performing the task.
In both Episode 46 (middle row) and Episode 49 (bottom row), confidence begins low and below the threshold. However, given the knowledge that calibration improves throughout the task horizon, we may prefer to allow the robot to proceed with the task as long as it is not contacting or nearly contacting any objects. Comparing these examples shows the potential of such an approach. In Episode 46, although confidence increases throughout the beginning of the task, it dives sharply after 40% task completion and remains low throughout the rest of the task, as the wine bottle is ultimately knocked down and the model is unable to recover. Given our proposed strategy, the task might have been halted before the wine bottle was contacted (at 40% completion), and could, for example, have been deferred to a human. On the other hand, in Episode 49 confidence estimates remain above the threshold for the rest of the trial, and the robot is successful. This episode also highlights the usefulness of an adaptive confidence threshold, given that the confidence estimate at 98% task completion would have fallen below the threshold at other timesteps.
More such qualitative examples are provided in Appendix Figure 18. These include cases where the strategy succeeds and others where it fails (e.g., confidence falls below the threshold while grasping in an ultimately successful trial).
4.5 Calibration Across Action Dimensions
Token-based VLAs such as OpenVLA or RT-2-X decompose a low-level command into 7 different action dimensions, i.e. Cartesian displacements, yaw-pitch-roll rotations, and a binary gripper toggle, predicted sequentially by an image-conditioned autoregressive LLM. Our baseline confidence estimate collapses this structure into a single scalar by averaging the top-token probabilities across the dimensions (Equation (3)); implicitly, this assumes that every dimension is calibrated to the same degree. That assumption may not hold in practice: a gripper ‘open’ or ‘close’ token may appear in nearly every demonstration, whereas a $90^{\circ}$ wrist roll could be rare. Such dataset imbalances and other implementation details might skew calibration across dimensions, and our relevant algorithms, e.g., for post hoc recalibration, might benefit from explicitly addressing this.
To study this question, our experiment consists of two parts:
  1. 1.
    Per-dimension calibration audit. For each dimension $d$ we treat $\max_{k}p^{(d)}_{t,k}$ as that dimension’s confidence and compute $\text{ECE}_{1}$, to examine whether different degrees of freedom are differentially calibrated.
  2. 2.
    Targeted recalibration. We compare classic Platt scaling to the action-wise variant of Section 3.6, which learns independent scaling parameters for each action dimension before averaging the transformed confidences across dimensions.
We perform our study using the baseline confidence estimates and all fine-tuned model variants (full, Quant-8, Quant-4) on the Spatial and Goal task suites. Each experiment is run for 1000 trials, with random 20%/80% calibration-test splits. Given the reduced size of the test set, we measure calibration with 10 equal-mass bins (instead of 12 as in other experiments). Beyond traditional Platt scaling, we also include a temperature scaling (Guo et al., 2017) baseline, where a single temperature is applied to all dimensions’ logits before the softmax. Figure 9 shows results with the baseline estimates for the full precision OpenVLA models; Appendix Figure 19 includes the results with the quantized models.
Two key observations arise:
  1. 1.
    $\text{ECE}_{1}$ varies by up to 5 times across dimensions, with no consistent best or worst dimensions across settings. It follows that a single scalar confidence masks significant differences in dimension-wise confidence estimates.
  2. 2.
    Replacing global Platt scaling with per-action-dimension transforms consistently lowers calibration error (and also improves over the temperature scaling baseline). ECE is improved by over 20% in some cases, without changing the selected tokens or adding meaningful runtime overhead.
Together, these results indicate that VLA calibration research should treat each degree of freedom as its own concern in order to improve effectiveness of relevant algorithmic tools. Dimension-aware post hoc methods such as action-wise Platt scaling offer a simple yet effective path toward that goal. More broadly, our findings underscore that calibrating VLAs introduces challenges not seen in other domains, so meaningful progress will demand substantial domain-specific research rather than wholesale adoption of understanding and techniques from other areas of deep learning.
Figure 9: On the left of each pair of plots, we compare miscalibration across action dimensions. On the right, we compare the performance of typical Platt scaling (and temperature scaling) to action-wise Platt scaling.
5 Discussion
Our work set out to answer a simple but previously unexplored question: can existing VLA policies tell us, in a statistically reliable way, how likely they are to succeed? Our experiments offer some preliminary answers. First, we find that better task performance correlates with lower calibration error, indicating that optimization for task performance and reliable confidence estimation are not in tension. Second, a zero-training algorithm using an ensemble of rephrased natural language instructions reduces expected calibration error by similar-to\!$\sim$20–40across 3 LIBERO suites and OpenVLA variants. This suggests that some miscalibration is simply high-variance noise in the language channel, and highlights a direction for future work in more reliable action prediction via perceptual perturbation. Third, our results indicate that miscalibration is not uniform across potential sources of confidence estimates. Instead, it varies over time and across action dimensions, and adaptive post hoc fixes (e.g., time-horizon-aware monitoring and action-wise Platt scaling) can exploit that structure. Together, our findings help establish a research agenda for VLA calibration: understanding how task accuracy and calibration co-evolve across diverse environments, leveraging natural perceptual variance via ensembles, incorporating temporal and per-dimension confidence dynamics, and developing rigorous algorithms that exploit these structures for identifying and addressing uncertainty.
6 Limitations
All experiments use OpenVLA fine-tuned in the LIBERO simulator. While this choice enables the controlled resets and hundreds of rollouts per task suite needed for stable calibration estimates, it limits architectural and domain diversity. However, running the same study on physical robots would require thousands of real executions, which can be prohibitively time-consuming and expensive. Simulation therefore may offer the most practical testbed for calibration research for the foreseeable future, although it cannot accurately model the sensor noise, latency, wear-and-tear, or other physical factors that may influence confidence calibration in the real world.
7 Future Work
Several extensions follow naturally from our work. Confidence calibration should be benchmarked across other environments, robot embodiments, and VLA architectures, including diffusion-based planners, continuous regression heads, and flow-matching controllers. Establishing strong baselines for these architectures will likely require new confidence surrogates and post hoc adjustments. Likewise, perturbation techniques besides instruction paraphrasing, e.g., inserting synthetic lighting changes or random distractors, could expose additional failure modes and inspire new multimodal ensemble techniques. Running such experiments on real robots will be essential for understanding how various physical factors affect VLA calibration in the real world. Finally, our temporal analysis hints at a period mid-trajectory where confidence is most trustworthy; integrating that signal into selective execution, planning, or human-in-the-loop systems is an open challenge for designing adaptive risk-aware robotic pipelines.
Overall, we hope that our work lays the groundwork for the rich body of research on calibration in VLAs necessary to place these models at the center of trustworthy real-world robotics systems.
Acknowledgements
We thank Yunzhu Li and Yixuan Wang for the helpful discussions and feedback. We also thank ONR Grant N00014-23-1-2436 for its generous support. This work is supported by the funds provided by the National Science Foundation and by DoD OUSD (R&E) under Cooperative Agreement PHY-2229929 (The NSF AI Institute for Artificial and Natural Intelligence).
Appendix A Additional Experiment Details
A.1 Rephrasing Prompts
The prompts given to GPT-4o-mini for rephrasing instructions from the LIBERO robot simulation environment are listed in Tables 5 and  6. Prompt 1 is the main prompt used throughout the experiments, and Prompts 2 and 3 are used to ablate the sensitivity to the rephrasings. Note that the different prompts lead to substantially different rephrasings, as Prompts 2 and 3 mandate the retention of the words “pick” and “place”, while as Table 1 shows, Prompt 1 leads to these words being replaced.
Prompt
Prompt 1 You are generating alternative phrasings of a robotic task instruction while preserving its exact meaning. ### Task Instruction: ’[TASK DESCRIPTION]’ ### Instructions: - Generate **20** alternative ways to phrase the task instruction. - Keep each instruction **concise and unambiguous**. - Ensure the instructions remain suitable for a **robot, not a human**. - Only make **semantically meaningless** changes (e.g., word order, synonyms, slight rewording). - Double-check that the new instructions mean the same exact thing for the robot; do not just substitute synonyms without considering context. - Do **not** introduce additional steps, remove essential details, or alter the action. ### Output Format: Each rephrased instruction should be wrapped in ‘[instruction]‘ and ‘[/instruction]‘ tags, like this: [instruction] Rephrased instruction 1 [/instruction] [instruction] Rephrased instruction 2 [/instruction]
Prompt 2 You are generating alternative phrasings of a robotic task instruction while preserving its exact meaning. ### Task Instruction: ’[TASK DESCRIPTION]’ ### Instructions: - Generate **20** alternative ways to phrase the task instruction. - Keep each instruction **concise and unambiguous**. - Ensure the instructions remain suitable for a **robot, not a human**. - Only make **semantically meaningless** changes (e.g., word order, synonyms, slight rewording). - Double-check that the new instructions mean the same exact thing for the robot; do not just substitute synonyms without considering context. - Do **not** introduce additional steps, remove essential details, or alter the action. - The first word of the instruction should be ‘PICK’, and then it should also include the word ‘PLACE’. ### Output Format: Each rephrased instruction should be wrapped in ‘[instruction]‘ and ‘[/instruction]‘ tags, like this: [instruction] Rephrased instruction 1 [/instruction] [instruction] Rephrased instruction 2 [/instruction]
Table 5: Prompts 1 and 2 for GPT-4o-mini to rephrase LIBERO task instructions. These prompts are used for Spatial and Object. There is a small modification for Goal.
Prompt
Prompt 3 You are generating alternative phrasings of a robotic task instruction while preserving its exact meaning. ### Task Instruction: ’[TASK DESCRIPTION]’ ### Instructions: - Generate **20** alternative ways to phrase the task instruction. - Keep each instruction **concise and unambiguous**. - Ensure the instructions remain suitable for a **robot, not a human**. - Only make **semantically meaningless** changes (e.g., word order, synonyms, slight rewording). - Double-check that the new instructions mean the same exact thing for the robot; do not just substitute synonyms without considering context. - Do **not** introduce additional steps, remove essential details, or alter the action. - Make the changes as minor as possible, as the robot’s language system is not very robust to rephrasing. - The first word of the instruction should be ‘PICK’, and then it should also include the word ‘PLACE’. ### Output Format: Each rephrased instruction should be wrapped in ‘[instruction]‘ and ‘[/instruction]‘ tags, like this: [instruction] Rephrased instruction 1 [/instruction] [instruction] Rephrased instruction 2 [/instruction]
Table 6: Prompt 3 for GPT-4o-mini to rephrase LIBERO task instructions. These prompts are used for Spatial and Object. There is a small modification for Goal.
Appendix B Additional Experiment Results
B.1 Relationship Between Task Success and Calibration
Figure 10 shows the relationship between calibration error and task error using the prompt ensemble method for estimating confidence.
Figure 10: Comparison of (4 different) calibration error measurements against task error rates for all model variants and task suites, using the Reprompt method to produce confidence estimates. A clear trend emerges: lower task error rates are associated with lower calibration error.
B.2 Prompt Ensembling
Figure 11 shows trends in percent change in calibration error using the prompt ensemble method. Figure 12 offers results for the ablation across ensemble sizes using the Goal and Object task suites.
Model Dataset Method $\text{ECE}_{1}$ $\text{ECE}_{2}$ Brier score NLL
Quant-4 Spatial Baseline 0.067 0.086 0.149 0.503
Reprompt 0.055 0.079 0.148 0.482
Object Baseline 0.091 0.112 0.135 0.489
Reprompt 0.066 0.094 0.132 0.454
Goal Baseline 0.161 0.182 0.231 0.748
Reprompt 0.146 0.154 0.224 0.672
Table 7: Calibration error measurements using the 4-bit fine-tuned model version and 3 different LIBERO task suites, for 2 different methods of confidence estimation: (1) baseline (2) ensembling over semantically equivalent prompts (“Reprompt”). The prompt ensemble method improves all measures across all settings.
Figure 11: Percent change in calibration error measurements by ensembling over semantically equivalent prompts.
Figure 12: Ablating the number of prompts in the prompt ensemble for the Goal and Object task suites. Increasing the number of prompts generally improves calibration error.
B.3 Calibration Across Task Time
Figures 13, 14, 15, 16, and 17 offer additional results on calibration across the task time horizon. Figure 18 shows further qualitative examples from the context-aware monitoring experiment.
Figure 13: Empirical study of calibration error across task time horizon for the Spatial task suite and the Quant-8 OpenVLA model.
Figure 14: Empirical study of calibration error across task time horizon for the Object task suite and the full precision OpenVLA model.
Figure 15: Empirical study of calibration error across task time horizon for the Object task suite and the Quant-8 OpenVLA model.
Figure 16: Empirical study of calibration error across task time horizon for the Goal task suite and the full precision OpenVLA model.
Figure 17: Empirical study of calibration error across task time horizon for the Goal task suite and the Quant-8 OpenVLA model.
Figure 18: Qualitative examples of a context-aware confidence calibration strategy for the task “put the wine bottle on the rack”. The red dashed line represents the 10% quantile of the confidence estimates output by the model across the task horizon, representing a potential threshold below which the robot may abstain from performing the task.
B.4 Calibration Across Action Dimensions
Figure 19 shows additional results for the action-wise recalibration experiments.
Figure 19: On the left of each pair of plots, we compare miscalibration across action dimensions. On the right, we compare the performance of typical Platt scaling (and temperature scaling) to action-wise Platt scaling.