### v2

# Confidence Calibration in Vision-Language-Action Models

Abstract
Trustworthy robot behavior requires not only high levels of task success but also that the robot can reliably quantify how likely it is to succeed. To this end, we present a first-of-its-kind study of confidence calibration in vision-language-action (VLA) foundation models, which map visual observations and natural language instructions to low-level robot motor commands. We establish a confidence baseline for VLAs, examine how task success relates to calibration error and how calibration evolves over time, and introduce two lightweight techniques to remedy the miscalibration we observe: prompt ensembles and action-wise Platt scaling. Our aim in this study is to begin to develop the tools and conceptual understanding necessary to render VLAs both highly performant and highly trustworthy via reliable uncertainty quantification.
1 Introduction
Confidence calibration, or the degree to which a model’s predicted probabilities reflect the true likelihood of its predictions being correct, is a cornerstone of reliable machine learning systems (guo2017calibration). When a model is well-calibrated, downstream decision-makers (e.g., humans, planning algorithms, or safety monitors) can trust that a confidence estimate of 95% implies that the predicted outcome will occur roughly 95% of the time (see Figure 1). Mismatches between confidence and actual outcomes can have severe negative consequences in high-stakes settings such as medical diagnosis or autonomous driving, motivating a rich literature on improving calibration in deep learning models (minderer2021revisiting; lakshminarayanan2017simple; gal2016dropout; tian2023just; lin2023generating; kadavath2022language).
Figure 1: To be trustworthy, a robotic system must be able to reliably express its confidence in its ability to perform a task, especially in high-stakes and open-world domains. A well-calibrated robot policy produces confidence estimates that align with its probability of task success. For example, the robot should succeed on 95% of instances for which it expresses 95% confidence.
Recently, the field of robotics has embraced a new class of vision-language-action (VLA) foundation models (brohan2023rt2visionlanguageactionmodelstransfer; kim2024openvlaopensourcevisionlanguageactionmodel; black2024pi0visionlanguageactionflowmodel; black2025pi05visionlanguageactionmodelopenworld). Leveraging large-scale, multistage pretraining, these models translate visual observations and natural language instructions into low-level joint-space commands, unifying multimodal perception with motor control. Although still a relatively new paradigm, current VLA systems already demonstrate previously unattainable generalization across environments, tasks, and robot embodiments. Because these models encode broad semantic and visuomotor priors from large-scale pretraining, they can be efficiently fine-tuned to new robots and downstream tasks, yielding large gains over training from scratch.
Since VLAs are designed for closed-loop interaction with the physical world, knowing when and how strongly to trust their actions is critical. For example, consider a robot performing a task in a safety-critical environment, or attempting to manipulate a valuable, fragile object. In these scenarios, if the policy can accurately express uncertainty, then costly or dangerous accidents can be avoided, such as by refining the instruction or deferring the task to a human. Despite the importance of calibrated confidence, basic questions remain largely unaddressed, e.g., whether VLAs are calibrated, how calibration evolves over the task horizon, and how it can be improved.
Contributions. To bridge this gap, we present (to our knowledge) the first study of calibration in VLAs, identifying key open questions and introducing practical remedies for miscalibration. Our contributions include:
  1. 1.
    Benchmarking task success vs. calibration. We evaluate the relationship between task success and calibration error across multiple benchmarks and VLA variants, finding that the model architecture and training objective may play significant roles in determining this relationship.
  2. 2.
    Prompt ensembles. We propose a lightweight, Bayesian-style method that averages a VLA’s confidence across multiple semantically equivalent rephrasings of an instruction. This approach consistently improves calibration, cutting expected calibration error by more than 20% on average.
  3. 3.
    Calibration over task time. We analyze calibration over task time, showing that confidence is often most reliable after making some task progress, suggesting natural points for risk-aware intervention.
  4. 4.
    Action-wise scaling. We discover systematic over-/under-confidence in different action dimensions and propose a method to recalibrate each action dimension independently to produce more reliable confidence estimates.
Our aim in this study is to begin to develop the tools and conceptual understanding necessary to render VLAs not only highly performant but also highly trustworthy via reliable uncertainty quantification.
2 Related Work
2.1 Confidence Calibration
A model is considered well-calibrated when the confidence (i.e., probability) it assigns to an outcome matches the long-run frequency of that outcome. Deviation from this condition is often quantified via expected calibration error (ECE) (guo2017calibration), which is approximated by grouping predictions with similar confidence into bins and measuring the absolute difference between confidence and accuracy in each bin. Beyond ECE, other metrics such as maximum calibration error (guo2017calibration), Brier score (brier1950verification), and negative log-likelihood (NLL) are used to capture related and complementary notions of miscalibration. Despite their impressive accuracy and trends toward better calibration, modern neural networks still display some persistent miscalibration (minderer2021revisiting), especially when the task distribution does not match the training distribution (ovadia2019trust). Consequently, a rich toolbox of post hoc fixes has emerged: Platt scaling (Platt1999), temperature scaling (guo2017calibration), histogram binning (zadrozny2001obtaining), and other recalibration methods (kumar2020verifieduncertaintycalibration; naeini2015obtaining; zollo2024improvingpredictorreliabilityselective) can be applied to an already trained model to reduce calibration error without altering its decision rule. Other popular methods to improve calibration of neural networks include dropout (gal2016dropout) and ensembling (lakshminarayanan2017simple).
2.2 Vision-Language-Action Models
Vision-Language-Action models take visual data and natural language instructions as input and output robot actions (brohan2023rt2visionlanguageactionmodelstransfer; kim2024openvlaopensourcevisionlanguageactionmodel; black2024pi0visionlanguageactionflowmodel; black2025pi05visionlanguageactionmodelopenworld). They are typically initialized from a visually-conditioned language model (VLM) that, in turn, is initialized from a pretrained large language model (LLM). This allows VLAs to leverage rich multimodal priors to integrate perception and action generation into a single end-to-end pipeline. While some VLAs retain the token-based output paradigm inherited from these base models (brohan2023rt2visionlanguageactionmodelstransfer; kim2024openvlaopensourcevisionlanguageactionmodel), others have augmented the architecture with, e.g., flow matching action experts for smooth, high-frequency control (black2024pi0visionlanguageactionflowmodel; black2025pi05visionlanguageactionmodelopenworld). These systems already perform complex, language-specified tasks across diverse environments and robot embodiments. However, they lack a reliable mechanism for quantifying the uncertainty of their chosen action sequences.
2.3 Calibration in LLMs
Given that most state-of-the-art VLAs are built on LLM backbones, it is natural to consider how our work relates to existing studies of calibration in LLMs, an area that has received considerable attention recently. Typical approaches to generating confidence estimates with LLMs include expressing calibration as a multiple-choice question for the LLM (kadavath2022language), directly verbalizing confidence with the model’s text output (lin2022teaching; tian2023just; band2024linguistic), or measuring semantic consistency across many sampled outputs for the same query (kuhn2023semantic; duan2024shifting; chen2024inside). However, it is difficult to adapt these methods to the VLA setting. For example, while LLMs may produce the entire sequence before measuring confidence, in robotics potential failures must be flagged much earlier in the trajectory to ensure safety and avoid costly accidents. Also, sampling multiple full trajectories in the physical world may be impossible. Finally, current VLAs lack the flexible and robust text-to-text interface of LLMs, and thus cannot be expected to, e.g., answer natural language questions about their confidence in an action prediction.
2.4 Uncertainty Quantification in Robotics
Uncertainty quantification has long been a focus in robotics, particularly through the lens of probabilistic robotics (thrun2005probabilistic; deisenroth2011pilco). Recent work explores the use of model ensembles in control systems, examining how disagreements among multiple models can capture uncertainty and guide more cautious policy updates and actions (chua2018deep; pathak2019selfsupervisedexplorationdisagreement; ataei2024dadeewellcalibrateduncertaintyquantification). Other advances include uncertainty-based active exploration (wang2023coplanner) and the application of conformal methods, which can provide distribution-free guarantees for planning or classification pipelines (ren2023robotsaskhelpuncertainty; sun2023conformal). How to obtain calibrated, task-level uncertainty estimates given the expressive, multimodal perception and control abilities of VLAs remains an open and safety-critical problem.
3 Calibration in Vision-Language-Action Models
In this section, we formalize the problem of confidence calibration for vision-language-action models. We define calibration, describe a typical example of how contemporary VLA architectures generate actions and how to extract a corresponding confidence estimate, and introduce the standard calibration metrics used in our empirical study (see Figure 2 for a visual summary of this material). Finally, we present two lightweight remedies for the miscalibration we observe in practice: prompt ensembles and action-wise Platt scaling.
3.1 Calibration
Let $C\!\in[0,1]$ denote the confidence reported by a robot policy and $Y\!\in\{0,1\}$ the binary indicator of task success (we use uppercase $C,Y$ for random variables and lowercase $c,y$ for their realizations). A perfectly calibrated predictor satisfies
$\mathbb{P}(Y=1\mid C=c)\;=\;c,\qquad\forall\,c\in[0,1].$ (1)
If this condition is met, then for the subset of trials on which the robot reports 80% confidence, we should observe successful completion 80% of the time.
Figure 2: Given an input image and text instruction, popular VLAs such as OpenVLA and RT-2 generate a distribution over discrete action tokens for each of the robot’s degrees of freedom. Confidence in each dimension’s prediction can be estimated using the probability assigned to the predicted token; a single estimate can be produced by averaging across dimensions. Given an uncalibrated confidence estimate, recalibration methods such as Platt scaling use a small calibration dataset to learn a map from uncalibrated confidence estimates to calibrated ones. Calibration error can be measured by comparing confidence estimates to actual task success rates.
3.2 Vision-Language-Action Models
At task timestep $t$, a VLA policy $\pi_{\theta}$ has access to the observed history $o_{t}=(v_{\leq t},l_{\leq t},a_{<t})$, where $v$ is a visual observation, $l$ is the (possibly fixed) natural language instruction, and $a$ is an action. The policy induces a distribution over the next action $\pi_{\theta}(a\mid o_{t}).$ When performing a task, the policy executes the most likely action $a_{t}^{\star}=\mathop{\rm argmax}_{a}\pi_{\theta}(a\mid o_{t})$. Given the need for calibration, we must extract a scalar confidence score $c_{t}$ from the model. We will interpret $c_{t}$ as an estimate of the probability that the task will ultimately succeed given the observation history and the chosen action, i.e., $\mathbb{P}(Y=1\mid o_{t},a_{t}^{\star})$.
Next, we propose a baseline method for extracting $c_{t}$ from a broad class of token-based VLAs. However, our subsequent analysis treats $c_{t}$ as a black-box number, so the following sections apply equally to token-based, diffusion-based, or other controllers, as long as they are able to emit such a scalar.
3.3 Baseline Confidence Estimation
Many state-of-the-art VLAs, including OpenVLA (kim2024openvlaopensourcevisionlanguageactionmodel) and RT-2 (brohan2023rt2visionlanguageactionmodelstransfer), represent an action with $D$ discrete tokens, where each represents one dimension of the robot’s action space. For each dimension $d\in\{1,\dots,D\}$ the policy outputs logits $z^{(d)}_{t}$ and probabilities
$p^{(d)}_{t}=\operatorname{softmax}\!\bigl(z^{(d)}_{t}\bigr),\qquad p^{(d)}_{t,k}=\pi_{\theta}\!\bigl(A^{(d)}_{t}=k\mid o_{t}\bigr)$ (2)
for action tokens $k\in\{1,\dots,K\}$ in an action vocabulary of size $K$. At time $t$, the policy selects the top tokens $a^{(d)}_{t}=\mathop{\rm argmax}_{k}p^{(d)}_{t,k}$ for each action dimension $d$, and decodes them into a continuous action for execution by the robot.
Because VLAs with token-based decoders closely mirror LLMs, we can adopt the usual LLM heuristic of using the probability of the token actually chosen as a confidence signal. For each action dimension, we take the probability assigned to the selected token, and average these values:
$c_{t}\;=\;\frac{1}{D}\sum_{d=1}^{D}\max_{k}p^{(d)}_{t,k}$. (3)
Averaging plays a role similar to length-normalization in LLMs (kuhn2023semantic), preventing confidence from being unfairly reduced for robots with many degrees of freedom.
3.4 Measuring Calibration
Calibration metrics translate deviations from the condition in Equation (1) into quantitative measures of miscalibration. Given the difficulty of measuring such a condition (i.e., comparing two distributions), we consider a range of metrics.
For what follows, let $\{(c_{i},y_{i})\}_{i=1}^{N}$ denote the reported confidence and binary outcome for each of $N$ robot trials (episodes). Each trial $i$ consists of timesteps $t=1,\ldots,T_{i}$ with per-timestep confidences $c_{i,t}$; to obtain a single trial-level value $c_{i}$ we apply an aggregation function $h$: $c_{i}=h(\{c_{i,t}\}_{t=1}^{T_{i}})$. Possible choices include the confidence before the first action, the mean across timesteps, or the min/max over the trajectory. The following measures are agnostic to choice of $h$; in our experiments we use the pre-action confidence $c_{i}=c_{i,1}$, reflecting the high-stakes open-world robotics setting in which early risk assessment is particularly valuable.
Expected Calibration Error.
One popular measure of miscalibration is expected calibration error (ECE) (guo2017calibration; fisch2022selective):
$\text{ECE}_{q}\;=\;\bigl(\mathbb{E}_{C}\!\bigl[\lvert\mathbb{P}(Y=1\mid C)-C\rvert^{q}\bigr]\bigr)^{\!1/q}\!.$ (4)
Put simply, ECE measures the expected difference between confidence and accuracy over the robot’s task data distribution. The parameter $q$ is typically set to $q\in\{1,2\}$, where $\text{ECE}_{1}$ weights deviations linearly and $\text{ECE}_{2}$ penalizes larger errors.
Because we can observe only a finite sample of trial results, the conditional expectation in Equation (4) cannot be directly measured. Instead, it is typically approximated with a binning-based estimator. With results from $N$ robot trials, we approximate the population ECE quantity in Equation (4) by first ordering predictions according to confidence, and splitting them into $M$ equal-sized bins $B_{1},\dots,B_{M}$. Then, our empirical estimate $\widehat{\text{ECE}}_{q}$ is given by
$\widehat{\text{ECE}}_{q}\;=\;\Bigl(\sum_{m=1}^{M}\frac{|B_{m}|}{N}\bigl|\mathrm{acc}(B_{m})-\mathrm{conf}(B_{m})\bigr|^{q}\Bigr)^{\!1/q},$ (5)
where
$\mathrm{acc}(B_{m})=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}y_{i},~~~~\mathrm{conf}(B_{m})=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}c_{i}.$
We report both $\text{ECE}_{1}$ and $\text{ECE}_{2}$ in our experiments.
Brier Score.
Brier score (brier1950verification) is a classic measure of the quality of probabilistic forecasting:
$\mathrm{BS}=\frac{1}{N}\sum_{i=1}^{N}(c_{i}-y_{i})^{2}.$ (6)
Brier score is an example of a proper scoring rule (Gneiting01032007), meaning that it is minimized only when the predicted distribution matches the true distribution for each example (as opposed to ECE, where data is grouped by confidence level). Brier score rewards both reliability (confidence matching success) and sharpness (predictions away from the population rate).
Negative Log-Likelihood.
Another metric used to measure calibration (guo2017calibration) is negative log-likelihood (NLL):
$\mathrm{NLL}=-\frac{1}{N}\sum_{i=1}^{N}\!\bigl[y_{i}\log c_{i}+(1-y_{i})\log(1-c_{i})\bigr],$ (7)
Also a proper scoring rule, NLL penalizes very confident failures much more heavily than Brier score.
Complementary Roles of Metrics.
Because each metric emphasizes a different facet of probabilistic quality, we report all three to obtain a detailed view of calibration performance. ECE focuses on reliability, or the alignment between stated confidence and empirical success, but is indifferent to how decisive those confidences are. A trivial predictor that outputs the base success rate on every trial achieves zero ECE while offering almost no utility for decision-making; Brier score and NLL would penalize this behavior. Conversely, since $Y$ is binary and only observed once, both Brier score and NLL implicitly discourage probabilistic predictions in the interior of $[0,1]$ when the goal is to achieve a very low loss. However, such estimates may sometimes be appropriate, e.g., due to randomness in the robot’s environment. Evaluating all three metrics together thus guards against corner cases and ensures that improvements in controlled evaluations translate into more actionable and reliable confidence estimates that are simultaneously calibrated, discriminative, and informative.
3.5 Prompt Ensembles
Similar to the VLMs from which they are derived (zhou2024analyzingmitigatingobjecthallucination), semantically meaningless lexical differences in instructions, e.g., “pick up the coffee cup” vs. “grab the mug”, might shift visual attention, alter path planning, and thereby change the predictions and confidence scores emitted by VLA models. Although such sensitivity can pose a challenge to reliable task execution, it also creates an opportunity to employ an ensemble-based approach (lakshminarayanan2017simple) to confidence estimation. Specifically, we can treat the particular wording of an instruction as a latent random variable and marginalize over it via Bayesian model averaging. Such averaging over rephrasings reduces variance in the final confidence estimate by canceling out the noise induced by word choice. Implementing this idea as an algorithm, we can then:
  1. 1.
    Generate rephrasings. An auxiliary LLM produces $r$ semantically equivalent prompts $\mathcal{L}_{alt}=\{l_{alt}^{(1)},\dots,l_{alt}^{(r)}\}$ (see Table 1 for examples).
  2. 2.
    Estimate confidence with each variant. Generating an action with $l_{alt}^{(i)}$ yields a confidence $c_{t}^{(i)}$ via Equation (3).
  3. 3.
    Aggregate. The final estimate is the ensemble mean $c^{ens}_{t}=\frac{1}{r}\sum_{i=1}^{r}c_{t}^{(i)}$.
Conceptually, this prompt ensemble technique can play a similar role to full model ensembles (lakshminarayanan2017simple) or inference-time dropout (gal2016dropout). With efficient batching on parallel hardware, this method can add little (i.e., roughly constant) wall-clock latency in practice.
Instruction Pick up the black bowl between the plate and the ramekin and place it on the plate
Rephrasings (1) Lift the black bowl located between the plate and the ramekin and set it on the plate.
(2) Grasp the black bowl found between the plate and the ramekin and move it to the plate.
(3) Take the black bowl positioned between the plate and the ramekin and position it on the plate.
Table 1: Example of multiple paraphrases of a single VLA instruction (from our experiments with LIBERO Spatial). The rephrasings carry the same meaning as the original instruction, enabling a reprompting approach in which we ensemble over predictions conditioned on lexically different yet semantically equivalent instructions.
3.6 Action-Wise Scaling
A standard remedy for miscalibration in classification models is to gather a small validation set from the task distribution and perform post hoc recalibration, learning a function that maps uncalibrated confidence estimates to calibrated ones (Platt1999; naeini2015obtaining; zadrozny2001obtaining; guo2017calibration). Post hoc recalibrators are typically simple functions of the original score function (e.g., logits) or confidence estimate.
One popular example of a post hoc recalibration method is Platt scaling (Platt1999; kumar2020verifieduncertaintycalibration). Given confidence outputs and binary outcomes $\{(c_{i},y_{i})\}_{i=1}^{N}$ on a held-out validation set, one fits an affine transform $g(c)=\sigma(\alpha c+\beta)$ that minimizes NLL,
$\min_{\alpha,\beta}-\sum_{i}\Bigl[y_{i}\log g(c_{i})+(1-y_{i})\log\!\bigl(1-g(c_{i})\bigr)\Bigr],$
where $\sigma(x)=1/(1+e^{-x})$. At inference, each $c_{t}$ is replaced with $\tilde{c}_{t}=g(c_{t})$, with the goal of aligning confidence with the probability of task success while leaving the model’s actions unchanged.
Unlike classification models that emit one predictive distribution per sample, many token-based VLAs output one predictive distribution per action dimension. Those dimensions can differ dramatically, for example because gripper open/close appears in nearly 100% of demonstrations, whereas “rotate wrist” is rarer. A single global transform therefore may not be able to correct all dimensions simultaneously.
Action-Wise Platt Scaling. We propose to address this heterogeneity by fitting one affine transform per dimension $d$:
$\tilde{c}_{t}\;=\;\frac{1}{D}\sum_{d=1}^{D}\sigma\!\bigl(\alpha_{d}\,\max_{k}p^{(d)}_{t,k}+\beta_{d}\bigr),$ (8)
with parameters $\{\alpha_{d},\beta_{d}\}$ learned on the same calibration set. Intuitively, $\alpha_{d}$ scales dimension $d$, flattening or sharpening its distribution, while $\beta_{d}$ shifts its overall optimism. As in standard Platt scaling, the chosen tokens $a^{(d)}_{t}$ remain unchanged; we only modify the reported confidence.
This dimension-wise perspective can be applied to other post hoc recalibrators as well, for instance action-wise temperature scaling. Beyond the specific methodology, we aim to highlight that VLA calibration will require domain-specific tools rather than a direct transplant of methods designed for standard classifiers.
4 Experiments
Having described the problem of confidence calibration in vision-language-action models, we next conduct an empirical investigation of the following questions:
  1. 1.
    How are task success and calibration error related, and what factors might play a role in this relationship? (Section 4.2)
  2. 2.
    Can prompt ensembles consistently improve confidence estimates? (Section 4.3)
  3. 3.
    How does calibration evolve over the task time horizon, and what does this suggest for potential monitoring systems? (Section 4.4)
  4. 4.
    Are action dimensions differentially calibrated, and what are the implications for recalibration? (Section 4.5)
Our aim is to highlight key issues and lay the empirical groundwork for future research on calibrating VLAs. [^1] Our code is available at https://github.com/thomaspzollo/vla_calibration.
4.1 Experiment Details
We perform our experiments using 4 different VLA variants: OpenVLA, MolmoAct (lee2025molmoactactionreasoningmodels), UniVLA (bu2025univlalearningacttaskcentric), and NORA (hung2025norasmallopensourcedgeneralist). We choose these models because they each predict tokens in a manner that allows for a natural probabilistic interpretation. [^2] Descriptions of confidence estimation with MolmoAct, UniVLA, and NORA are found in Appendix A.1. All models are fine-tuned on 4 different task suites from the LIBERO (liu2023liberobenchmarkingknowledgetransfer) benchmark: Spatial, Object, Goal, and 10. LIBERO is a simulation environment for language-conditioned robot manipulation tasks inspired by human activities (see Table 3 for task examples). We also include results for the 8-bit and 4-bit quantized versions of OpenVLA fine-tuned on Spatial, Object, and Goal, for a total of 22 model/task suite combinations.
Each task suite features 10 different tasks with 50 randomized initializations, for a total of 500 examples. For a given task suite, task success rate represents the proportion of the 500 trials that result in success (while task error rate represents the proportion of trials that result in failure). To calculate ECE, we use 12 equal-mass bins and the Python package released with kumar2020verifieduncertaintycalibration.
Unless otherwise noted, we focus on confidence estimates produced at the first timestep, before any action is taken. This aligns with the need for safety in open-world robot deployments, where robots should signal uncertainty as early as possible in order to avoid costly or dangerous incidents. In Section 4.4, we explore how calibration differs across timesteps over the task horizon.
4.2 Relationship Between Calibration and Task Success
Much of the early research on calibration in deep learning models was based on perceptions of how task success (usually image classification accuracy) was related to calibration error. In particular, influential work argued that modern neural networks are poorly calibrated (guo2017calibration), suggesting that improved accuracy comes at the expense of calibration. This led researchers to propose training interventions to augment the cross-entropy loss with other objectives seeking improved calibration, potentially at a cost to overall accuracy (pmlr-v80-kumar18a; mukhoti2020calibratingdeepneuralnetworks). However, subsequent research found that, in fact, more accurate networks are generally better calibrated (and easier to recalibrate), and thus modifications to training procedures may not be needed (minderer2021revisiting). Instead, techniques such as post hoc recalibration (guo2017calibration; zadrozny2001obtaining; kumar2020verifieduncertaintycalibration) and ensembling (lakshminarayanan2017simple; fort2024ensembleeverywheremultiscaleaggregation), applied to models trained for high accuracy, are sufficient to achieve low calibration error (although calibration under distribution shift remains a significant challenge (ovadia2019trust)).
To establish high-level direction for calibration research in VLAs, our first experiment focuses on this important question of how task success relates to calibration error (according to $\text{ECE}_{1}$, $\text{ECE}_{2}$, Brier score, and NLL) across the 22 model/task suite combinations described above. Results are visualized in Figure 3, and also reported in Table 6. All models exhibit a roughly monotonic relationship between task error and the discriminative metrics (Brier score and NLL), but they differ on ECE: OpenVLA and MolmoAct tend to achieve lower ECE when task error is low, whereas UniVLA and NORA do not show as clean a trend. One possible explanation is that in OpenVLA and MolmoAct, the cross-entropy loss (a proper scoring rule) more directly supervises discretized action dimension tokens, so bin-averaged confidence better tracks per-task success. By contrast, the latent or compressed action representations and auxiliary objectives in UniVLA and NORA may introduce a more significant mismatch between token probabilities and success. These patterns are only suggestive, but they point to architectural complexity in modern VLAs as a potential source of unfavorable calibration behavior. Future work with controlled ablations is needed to fully understand which design choices drive the observed behavior.
Figure 3: Visualization of task error rates compared against 4 different calibration error measurements for 4 VLA variants (OpenVLA, MolmoAct, UniVLA, and NORA) and 4 LIBERO task suites (Spatial, Object, Goal, 10), as well as OpenVLA 8- and 4-bit versions on Spatial, Object, and Goal. All models exhibit a roughly monotonic relationship between task error and the discriminative measures (Brier score and NLL). ECE shows differences between model families, potentially due to architecture and objective differences.
4.3 Ensembling Confidence Across Prompts
We next study the empirical effectiveness of the prompt ensemble approach described in Section 3.5. First, for the natural language instruction associated with each task, we create 20 rephrasings using GPT-4o-mini (see Appendix Table 4 for the prompts used). During testing, we produce a confidence estimate conditioned on each “reprompt”, and average over these confidence scores to obtain the final model confidence. We evaluate OpenVLA and its quantized variants across 3 task suites, as well as UniVLA and NORA on the Spatial suite. We measure $\text{ECE}_{1}$, $\text{ECE}_{2}$, Brier score, and NLL, comparing to the baseline method for producing confidence estimates using the original instruction.
Detailed results are recorded in Table 2 (results for the 8-bit and 4-bit quantized models are deferred to Appendix Table 7). Across all models and task suites, the Reprompt (prompt ensemble) approach always decreases calibration error according to $\text{ECE}_{1}$, $\text{ECE}_{2}$, and NLL, and never performs worse according to any of the metrics. Decreases in ECE can reach up to 40%, and average reductions are roughly 20% (see Figure 7 for a visualization of percent changes across all metrics). The improvement is larger for ECE than the proper scoring rules (Brier score and NLL), where average miscalibration decreases are closer to 5-10%. This suggests that improvements in reliability (i.e., confidence matching marginal success rates) are generally greater than those in sharpness, which is expected from an ensemble technique targeted at reducing variance in confidence estimation. Overall, given its lightweight nature and strong empirical effectiveness, such data augmentation approaches seem like a promising direction for enhancing uncertainty quantification in VLAs.
Model Dataset Method $\text{ECE}_{1}$ $\text{ECE}_{2}$ Brier NLL
OpenVLA Spatial Baseline 0.088 0.106 0.150 0.533
Spatial Reprompt 0.052 0.068 0.145 0.477
Object Baseline 0.060 0.073 0.108 0.401
Object Reprompt 0.036 0.053 0.105 0.361
Goal Baseline 0.151 0.170 0.207 0.707
Goal Reprompt 0.115 0.132 0.197 0.620
UniVLA Spatial Baseline 0.162 0.193 0.064 0.282
Spatial Reprompt 0.157 0.192 0.064 0.262
NORA Spatial Baseline 0.099 0.115 0.105 0.386
Spatial Reprompt 0.092 0.112 0.105 0.376
Table 2: Calibration error measurements for 2 different methods of confidence estimation: (1) baseline (average selected token probability); (2) Reprompt (ensembling over semantically equivalent prompts). Prompt ensembling consistently improves all measures, and never leads to worse calibration.
To understand the robustness of these results, we perform multiple ablations; because of space considerations, we defer full details and results to Appendix B.3. First, we consider the effect of changing the prompt given to GPT-4o-mini for producing the 20 instruction rephrasings, finding that improvements in calibration error from the Reprompt method are robust to different rephrasing prompts. Second, we consider the effect of the number of prompts used in the prompt ensemble, and see that the algorithm behaves favorably, where calibration error generally decreases as more instructions are included in the ensemble.
4.4 Calibration Over Task Time
The previous experiments evaluate confidence before the first action is executed, a conservative choice for safety-critical deployments. Yet many tasks might allow the robot to collect more information without risk before having to express confidence. For instance, in the wine glass scenario of Figure 1, the gripper can hover above the stem, refine its scene representation, and only then decide whether it is confident enough to proceed. More context should, in principle, yield better calibration.
To test this intuition, we measure calibration across 500 test trials for 100 different levels of task completion ($\{0,1,\dots,99\}\%$), where task completion is calculated as the current timestep index $t$ divided by the total number of timesteps in the task episode (and multiplied by 100). Results for each level of task completion are averaged across the 500 trials, to study whether high-level confidence and calibration trends might occur. In this section, we focus on 6 model/task suite combinations (OpenVLA and its 8-bit version, each applied to Spatial, Object, and Goal suites), to understand how any observations might generalize. Our goal is to examine whether the quality of confidence estimates changes as the robot progresses in its task. At each measured timestep, we compute $\text{ECE}_{1}$ and Brier score using the baseline method for producing confidence estimates. Because a downstream safety monitor might consider the history of recent estimates (beyond the current one), we evaluate three aggregation rules for reporting a confidence estimate at each timestep:
  1. 1.
    Current - confidence from the current timestep only;
  2. 2.
    Window (5) - mean over the baseline confidence estimates from the current timestep and the four immediately preceding ones;
  3. 3.
    Avg. All - mean over the baseline confidence estimates from all steps seen so far.
Beyond calibration error, we also visualize how average confidence estimates evolve over the task horizon, separated by successful and failed trials. Finally, we plot reliability diagrams for the timesteps corresponding to $\{0,50,75,99\}\%$ completion. Reliability diagrams (guo2017calibration) offer a visualization of expected calibration error, using a similar binning strategy. They show confidence against accuracy for each bin, where a well-calibrated system lies on the $x=y$ line. Results for the Spatial task suite with the fine-tuned OpenVLA model are shown in Figure 4. Additional results for Spatial with the Quant-8 model (Figure 10), as well as the Object and Goal suites with the full precision (Figures 11, 13) and Quant-8 models (Figures 12, 14), are presented in Appendix B.4.
Figure 4: Empirical study of calibration error across the task time horizon. In the top row (a), the left two plots show how calibration evolves with task progress, while the right two plots show the average confidence by task time, grouped by successful and failed trials. The bottom two rows (b, c) offer a sample of the reliability diagrams produced by different methods for aggregating confidence estimates across time. Overall, these results illustrate that calibration can improve as the task progresses and more information is gathered, suggesting opportunities for context-aware uncertainty interventions.
Focusing first on results using the confidence estimate from the current step, a clear pattern emerges. As shown in Figure 4 and Figures 10-14, across all 6 settings and both metrics calibration improves sharply from 0% to approximately 50% completion, then plateaus or deteriorates back towards the original level (left-most and left-middle plots in section (a) of the results figures). Beyond lower scores on calibration error, our additional plots characterize the improved probabilistic nature of the confidence estimates towards the middle of the task horizon. In the right-middle and right-most plots in the top section (a) (titled “Successful Trials” and “Failed Trials”), the gap between the average confidence on successful vs. failed trials tends to be greatest around 50% task completion. Additionally, considering the top row of reliability diagrams (section (b)), we can see that the confidence estimates using the current step become far more reliable around this point: error for many bins is large at the beginning, then the difference between confidence and accuracy for most bins becomes smaller until the task is roughly 75% complete. Near task completion, the gap between confidence on successful vs. failed trials closes again, and reliability suffers as a result. These results suggest a practical recipe: let the robot execute a certified-safe prefix of the trajectory, and then assess its confidence and intervene if necessary. Such horizon-aware monitoring balances safety (no contact made during the prefix) with the improved reliability in confidence estimates that comes from extra perceptual context.
We next observe the effects of other simple strategies for aggregating confidence estimates across timesteps. As described above, alongside the current step confidence, we also consider a sliding window average over 5 single-step estimates and the average of all observed single-step estimates. Across all model/task suite combinations, the sliding window method is particularly effective at improving the Brier score around the middle of the task horizon, although ECE is not always improved. With respect to averaging all single-step estimates so far, we observe that Brier score improves throughout the task, but ECE generally gets worse. The underlying behavior driving these changes can be observed in the bottom row of reliability diagrams (marked (c)): the “Avg. All” method is successful at assigning relatively lower confidence to failed examples, but also reduces variance in the estimates such that they are highly overconfident in these bins (i.e., for the leftmost bin at 99% completion, confidence is still high, but accuracy is near zero). Thus, while this approach fares poorly according to ECE, it is actually promising in the sense that it enables more effective discrimination between successful and failed trials.
Finally, to further probe the generality of these results, we repeat these experiments using the UniVLA model, presented in Appendix Figures 15, 16, and 17. We once again find that calibration improves after making some task progress before deteriorating again, suggesting opportunities for context- and risk-aware applications of confidence quantification.
4.4.1 Qualitative Examples of Context-Aware Confidence Monitoring
To illustrate how context-aware confidence monitoring could work in practice, we apply it to a representative pick and place task from the Goal suite. The task is to “put the wine bottle on the rack”, a case where the robot system should be relatively conservative to avoid breaking glass. We consider a naive approach to context-aware monitoring, proposing to halt task performance when both: (1) the confidence level falls below a threshold set to the 10% quantile of confidence estimates for that point in the task horizon across all task trials (based on percent completion); (2) the robot is within a few inches of contacting an object, or already has contacted an object. Since building a system to detect proximity to objects is beyond the scope of this work, we focus on a qualitative demonstration of this idea. For each example under examination, we plot current confidence for $\{0,20,40,60,80,98\}\%$ completion, as well as the corresponding 10% quantile risk threshold and an image of the robot environment at that time.
Figure 5: Qualitative examples of a context-aware confidence monitoring strategy applied to a task from the Goal suite. Here, the task is to “put the wine bottle on the rack”. The red dashed line represents the 10% quantile of the confidence estimates output by the model across the task time horizon, offering a potential threshold below which the robot may abstain from performing the task.
Some particularly illustrative examples are shown in Figure 5. First, we consider the episode shown across the top row (marked (a), Episode 29 of the 50 in the Goal task suite). Here, confidence begins high, but falls until it is slightly below the halting threshold at 20% task completion, coinciding with the difficult subtask of gripping the rounded glass bottle. Allowed to continue, the model is able to grip the bottle, and confidence rebounds as it approaches the wine rack. However, confidence once again falls quickly when it begins to set the bottle down, possibly because the difficult grip put the bottle in an unfavorable position for placement. In the end, the robot fails to place the bottle securely on the rack, with potentially negative real-world consequences. Given the opportunity, a safety intervention could have been performed at multiple points before this incident, either to avoid grasping the bottle in the first place or to reset the bottle safely on the table instead of trying a failed placement.
In both Episode 46 (marked (b)) and Episode 49 (marked (c)), confidence begins below the threshold. However, given the knowledge that calibration improves throughout the task horizon, we may prefer to allow the robot to proceed with the task as long as it is not too near any objects. Comparing these examples shows the potential of such an approach. In Episode 46, although confidence increases throughout the beginning of the task, it dives sharply after 40% task completion and remains low throughout the rest of the task, as the wine bottle is ultimately knocked down and the model is unable to recover. Given our proposed strategy, the task might have been halted before the wine bottle was contacted (at 40% completion), and could, for example, have been deferred to a human. On the other hand, in Episode 49 confidence estimates remain above the threshold for the rest of the trial, and the robot is successful. This episode also highlights the usefulness of an adaptive confidence threshold, given that the confidence estimate at 98% task completion would have fallen below the threshold at other timesteps.
More such qualitative examples are provided in Appendix Figure 18. These include both cases where the strategy succeeds and others where it fails (e.g., in Episode 1 confidence falls below the threshold while grasping in an ultimately successful trial).
4.5 Calibration Across Action Dimensions
Some token-based VLAs such as OpenVLA or MolmoAct decompose a low-level command into tokens explicitly corresponding to their 7 different action dimensions. Our baseline confidence estimate collapses this structure into a single scalar by averaging the top-token probabilities across the dimensions; implicitly, this assumes that every dimension is calibrated to the same degree. That assumption may not hold in practice: a gripper ‘open’ or ‘close’ token may appear in nearly every demonstration, whereas a $90^{\circ}$ wrist roll could be rare. Such dataset imbalances and other implementation details might skew calibration across dimensions, and relevant algorithms, e.g., for post hoc recalibration, might benefit from explicitly addressing this. To study this question, our final experiment consists of two parts. (1) Per-dimension calibration audit. For each dimension $d$ we treat $\max_{k}p^{(d)}_{t,k}$ as that dimension’s confidence and compute $\text{ECE}_{1}$, to examine whether different degrees of freedom are differentially calibrated. (2) Targeted recalibration test. We compare classic Platt scaling to the action-wise variant of Section 3.6, which learns independent scaling parameters for each action dimension before averaging the transformed confidences across dimensions.
We perform our study using MolmoAct and all 3 OpenVLA variants (full, Quant-8, Quant-4) fine-tuned on the Spatial and Goal task suites. Each experiment is run for 1000 trials, with random 20%/80% calibration-test splits. Given the reduced size of the test set, we measure calibration with 10 equal-mass bins (instead of 12 as in other experiments). Beyond traditional Platt scaling, we also include a range of baselines: temperature scaling (guo2017calibration), histogram binning (zadrozny2001obtaining), and Platt binning (kumar2020verifieduncertaintycalibration). (Note: temperature scaling is excluded from MolmoAct results, as the variable-length action encodings preclude a straightforward adaptation.) Figure 6 shows results with the baseline estimates for the full precision OpenVLA and MolmoAct models; Appendix Figure 19 includes the results with the quantized models.
Two key observations arise. First, $\text{ECE}_{1}$ varies by up to 2 times across dimensions, with no consistent best or worst dimensions across settings. It follows that a single scalar confidence masks significant differences in dimension-wise confidence estimates. Second, we can observe that replacing global Platt scaling with per-action-dimension transforms consistently lowers calibration error (and also improves over the additional baselines). ECE is improved by over 20% in some cases, without changing the selected tokens or adding meaningful runtime overhead. Together, these results indicate that VLA calibration research should treat each degree of freedom as its own concern in order to improve the effectiveness of relevant algorithmic tools. Dimension-aware post hoc methods such as action-wise Platt scaling offer a simple yet effective path toward that goal. More broadly, our findings underscore that calibrating VLAs introduces challenges not seen in other domains, so meaningful progress will demand substantial domain-specific research rather than wholesale adoption of understanding and techniques from other areas of deep learning.
Figure 6: On the left of each pair of plots, we compare miscalibration across action dimensions. On the right, we compare the performance of baselines to action-wise Platt scaling.
5 Discussion
In this work, we set out to answer a simple but previously unexplored question: can existing VLA models tell us, in a statistically reliable way, how likely they are to succeed? Our experiments offer some preliminary answers. First, we find that average top-token probability is a reasonable confidence baseline, and that decisions like architecture and learning objectives may affect the crucial relationship between task performance and calibration error. Second, a zero-training algorithm using an ensemble of rephrased natural language instructions reliably reduces expected calibration error across 3 LIBERO suites and 3 VLA variants. This suggests that some miscalibration is simply high-variance noise in the language channel, and highlights a direction for future work in more reliable action prediction via input perturbation. Third, our results indicate that miscalibration is not uniform across potential sources of confidence estimates. Instead, it varies over time and across action dimensions, and adaptive post hoc fixes (e.g., time-horizon-aware monitoring and action-wise Platt scaling) can exploit that structure. Together, our findings point towards some interesting initial directions of study for VLA calibration: understanding how task accuracy and calibration co-evolve across diverse modeling approaches and environments, creating calibration tools grounded in practical VLA implementation details, incorporating temporal confidence dynamics, and developing rigorous algorithms that exploit these structures for identifying and addressing uncertainty.
6 Limitations
All experiments in this work are conducted in simulation. This choice is common in robotics research (e.g., lillicrap2015continuous; schulman2017proximal), and enables the controlled resets and hundreds of rollouts per task suite needed for stable calibration estimates. However, it limits domain diversity and leaves open questions of how these results translate to real settings. Replicating the same study on physical robot hardware would require thousands of physical executions, which can be prohibitively time-consuming and expensive. Accordingly, simulation remains the most practical testbed for calibration research in the near term, though it cannot fully capture sensor noise, latency, wear-and-tear, and other physical factors that may influence confidence calibration in the real world. Studying these questions and techniques on real robots will ultimately be essential for understanding how these and other factors affect VLA calibration under realistic conditions.
7 Future Work
Several extensions follow naturally from our work. Confidence calibration should be benchmarked across other environments, robot embodiments, and VLA architectures, including diffusion-based planners, continuous regression heads, and flow-matching controllers. Establishing strong baselines for these architectures will likely require new confidence surrogates and post hoc adjustments. Likewise, perturbation techniques besides instruction paraphrasing, e.g., inserting synthetic lighting changes or random visual distractors, could expose additional failure modes and inspire new multimodal ensemble techniques. Running such experiments on real robots will be essential for understanding how various physical factors affect VLA calibration in the real world. Uncertainty estimation also holds the potential to guide efficient data gathering and fine-tuning, as in active learning (Wang_2017). Further, while all of our experiments focused on fine-tuned in-distribution settings, future work should consider these phenomena in the zero-shot and out-of-distribution settings. Finally, our temporal analysis hints at a period mid-trajectory where confidence is most trustworthy; integrating that signal into selective execution, planning, or human-in-the-loop systems is an open challenge for designing adaptive risk-aware robotic pipelines.
Overall, we hope our work helps build a foundation for the research on VLA calibration needed to make these models trustworthy components of real-world robotic systems.
Acknowledgements
We thank Yunzhu Li and Yixuan Wang for the helpful discussions and feedback. We also thank ONR Grant N00014-23-1-2436 for its generous support. This work is supported by the funds provided by the National Science Foundation and by DoD OUSD (R&E) under Cooperative Agreement PHY-2229929 (The NSF AI Institute for Artificial and Natural Intelligence).
Appendix
Appendix A Additional Experiment Details
Task examples from the LIBERO task suites are shown in Table 3.
Dataset Example Image Example Instruction
Spatial [Uncaptioned image] Pick up the black bowl between the plate and the ramekin and place it on the plate.
Object [Uncaptioned image] Pick up the alphabet soup and place it in the basket.
Goal [Uncaptioned image] Open the middle drawer of the cabinet.
10 [Uncaptioned image] Put both the alphabet soup and the tomato sauce in the basket.
Table 3: Task examples from the LIBERO task suites: Spatial, Object, Goal, and 10.
A.1 Confidence Estimation with VLA Variants
MolmoAct
At each timestep, MolmoAct predicts a sequence of discrete action tokens for every action dimension, with a softmax distribution over the 256 action bins for each token. For a given action dimension, we take the probability assigned to the executed bin for each of its tokens and average these probabilities to obtain a per-dimension confidence. This yields one scalar confidence value per action dimension. We then average these per-dimension confidences across all action dimensions to obtain a single scalar confidence for the action at that timestep.
UniVLA
At each timestep, UniVLA predicts a fixed number (4) of discrete latent action tokens, each chosen from a codebook of size 16. For each latent token, we take the softmax probability assigned to the selected code as its token-level confidence. Because these latent tokens are jointly decoded into all continuous control dimensions, the confidence scores do not correspond to particular action dimensions. Finally, we average these per-token confidences to obtain a single scalar confidence for UniVLA at that timestep.
NORA
At each timestep, NORA predicts an action chunk encoded as a sequence of discrete tokens, each selected via a softmax over the augmented action vocabulary. For every token, we take the softmax probability of the sampled token as its token-level confidence. Because the NORA tokenizer mixes information from all control dimensions, we do not derive per-dimension confidences. Instead, we compute a single scalar confidence for NORA by averaging these token-level confidences across all tokens in the predicted action chunk for that timestep.
A.2 Rephrasing Prompts
The prompts given to GPT-4o-mini for rephrasing instructions from the LIBERO robot simulation environment are listed in Tables 4 and 5. Prompt 1 is the main prompt used throughout the experiments, and Prompts 2 and 3 are used to ablate the sensitivity to the rephrasings. Note that the different prompts lead to substantially different rephrasings, as Prompts 2 and 3 mandate the retention of the words “pick” and “place”, while as Table 1 shows, Prompt 1 leads to these words being replaced.
Prompt
Prompt 1 You are generating alternative phrasings of a robotic task instruction while preserving its exact meaning. ### Task Instruction: ’[TASK DESCRIPTION]’ ### Instructions: - Generate **20** alternative ways to phrase the task instruction. - Keep each instruction **concise and unambiguous**. - Ensure the instructions remain suitable for a **robot, not a human**. - Only make **semantically meaningless** changes (e.g., word order, synonyms, slight rewording). - Double-check that the new instructions mean the same exact thing for the robot; do not just substitute synonyms without considering context. - Do **not** introduce additional steps, remove essential details, or alter the action. ### Output Format: Each rephrased instruction should be wrapped in ‘[instruction]‘ and ‘[/instruction]‘ tags, like this: [instruction] Rephrased instruction 1 [/instruction] [instruction] Rephrased instruction 2 [/instruction]
Prompt 2 You are generating alternative phrasings of a robotic task instruction while preserving its exact meaning. ### Task Instruction: ’[TASK DESCRIPTION]’ ### Instructions: - Generate **20** alternative ways to phrase the task instruction. - Keep each instruction **concise and unambiguous**. - Ensure the instructions remain suitable for a **robot, not a human**. - Only make **semantically meaningless** changes (e.g., word order, synonyms, slight rewording). - Double-check that the new instructions mean the same exact thing for the robot; do not just substitute synonyms without considering context. - Do **not** introduce additional steps, remove essential details, or alter the action. - The first word of the instruction should be ‘PICK’, and then it should also include the word ‘PLACE’. ### Output Format: Each rephrased instruction should be wrapped in ‘[instruction]‘ and ‘[/instruction]‘ tags, like this: [instruction] Rephrased instruction 1 [/instruction] [instruction] Rephrased instruction 2 [/instruction]
Table 4: Prompts 1 and 2 for GPT-4o-mini to rephrase LIBERO task instructions. These prompts are used for Spatial and Object. There is a small modification for Goal.
Prompt
Prompt 3 You are generating alternative phrasings of a robotic task instruction while preserving its exact meaning. ### Task Instruction: ’[TASK DESCRIPTION]’ ### Instructions: - Generate **20** alternative ways to phrase the task instruction. - Keep each instruction **concise and unambiguous**. - Ensure the instructions remain suitable for a **robot, not a human**. - Only make **semantically meaningless** changes (e.g., word order, synonyms, slight rewording). - Double-check that the new instructions mean the same exact thing for the robot; do not just substitute synonyms without considering context. - Do **not** introduce additional steps, remove essential details, or alter the action. - Make the changes as minor as possible, as the robot’s language system is not very robust to rephrasing. - The first word of the instruction should be ‘PICK’, and then it should also include the word ‘PLACE’. ### Output Format: Each rephrased instruction should be wrapped in ‘[instruction]‘ and ‘[/instruction]‘ tags, like this: [instruction] Rephrased instruction 1 [/instruction] [instruction] Rephrased instruction 2 [/instruction]
Table 5: Prompt 3 for GPT-4o-mini to rephrase LIBERO task instructions. These prompts are used for Spatial and Object. There is a small modification for Goal.
Appendix B Additional Experiment Results
All task success and calibration results for 22 VLA/task suite combinations are shown in Table 6.
B.1 Relationship Between Task Success and Calibration
Table 6 features all results plotted in Figure 3.
Dataset Model Quant $\text{ECE}_{1}$ $\text{ECE}_{2}$ Brier NLL Succ. Rate
Spatial NORA - 0.099 0.115 0.105 0.386 0.894
Object NORA - 0.210 0.229 0.090 0.343 0.960
Goal NORA - 0.065 0.077 0.111 0.391 0.880
10 NORA - 0.184 0.212 0.260 0.811 0.674
Spatial MolmoAct - 0.113 0.123 0.134 0.625 0.860
Object MolmoAct - 0.046 0.059 0.062 0.296 0.936
Goal MolmoAct - 0.129 0.141 0.146 0.657 0.846
10 MolmoAct - 0.175 0.184 0.205 0.716 0.772
Spatial UniVLA - 0.162 0.193 0.064 0.282 0.972
Object UniVLA - 0.082 0.135 0.054 0.224 0.962
Goal UniVLA - 0.116 0.163 0.066 0.261 0.958
10 UniVLA - 0.134 0.174 0.092 0.339 0.932
Spatial OpenVLA - 0.088 0.106 0.150 0.533 0.828
Object OpenVLA - 0.060 0.073 0.108 0.401 0.880
Goal OpenVLA - 0.151 0.170 0.207 0.707 0.758
10 OpenVLA - 0.381 0.390 0.398 1.382 0.532
Spatial OpenVLA Quant-8 0.070 0.093 0.138 0.488 0.844
Object OpenVLA Quant-8 0.057 0.065 0.110 0.409 0.876
Goal OpenVLA Quant-8 0.140 0.163 0.205 0.713 0.760
Spatial OpenVLA Quant-4 0.067 0.086 0.149 0.503 0.822
Object OpenVLA Quant-4 0.091 0.112 0.135 0.489 0.850
Goal OpenVLA Quant-4 0.161 0.182 0.231 0.748 0.712
Table 6: All task success and calibration results for 22 VLA/task suite combinations.
B.2 Prompt Ensembling
Table 7 has prompt ensemble results for the quantized OpenVLA models. Figure 7 shows trends in percent change in calibration error using the prompt ensemble method.
Model Dataset Method $\text{ECE}_{1}$ $\text{ECE}_{2}$ Brier score NLL
OpenVLA Quant-8 Spatial Baseline 0.070 0.093 0.138 0.488
Reprompt 0.050 0.062 0.131 0.434
Object Baseline 0.057 0.065 0.110 0.409
Reprompt 0.041 0.052 0.108 0.375
Goal Baseline 0.140 0.163 0.205 0.713
Reprompt 0.117 0.152 0.194 0.608
OpenVLA Quant-4 Spatial Baseline 0.067 0.086 0.149 0.503
Reprompt 0.055 0.079 0.148 0.482
Object Baseline 0.091 0.112 0.135 0.489
Reprompt 0.066 0.094 0.132 0.454
Goal Baseline 0.161 0.182 0.231 0.748
Reprompt 0.146 0.154 0.224 0.672
Table 7: Calibration error measurements using the 8-bit and 4-bit fine-tuned model versions and 3 different LIBERO task suites, for 2 different methods of confidence estimation: (1) baseline (2) ensembling over semantically equivalent prompts (“Reprompt”). The prompt ensemble method improves all measures across all settings.
Figure 7: Percent change in calibration error measurements by ensembling over semantically equivalent prompts.
Figure 8: Ablation results for the prompt ensemble method, where different prompts are given to GPT-4o-mini for producing the rephrasings of the original instruction. Prompt 1 is used in the main experiments, while Prompts 2 and 3 are variants. Results show that improvements in calibration error from Reprompt are robust to different rephrasing prompts.
Figure 9: Ablating the number of prompts in the prompt ensemble for the Spatial, Goal, and Object task suites. Increasing the number of prompts generally improves calibration error.
B.3 Ablations for Prompt Ensembles
To understand the robustness of our prompt ensemble results, we perform multiple ablations using the OpenVLA model.
First, we consider the effect of changing the prompt given to GPT-4o-mini for producing the 20 instruction rephrasings. In addition to the prompt used to produce the results in Table 2, we try two additional prompts, recorded in Appendix Table 4 and Table 5; “Prompt 1” is the original prompt, and “Prompt 2” and “Prompt 3” represent new prompts for the ablation. Note that the different prompts lead to substantially different rephrasings, as Prompts 2 and 3 mandate the retention of the words “pick” and “place”, while as Table 1 shows, Prompt 1 leads to these words being replaced. The ablation is run across all 3 task suites using fine-tuned OpenVLA, and we measure $\text{ECE}_{1}$ and NLL. Results are shown in Figure 8. Improvements in calibration error from the Reprompt method are robust to different rephrasing prompts, with all ensemble variants producing lower error than the baseline method across both metrics and all 3 task suites.
Second, we consider the effect of the number of prompts used in the prompt ensemble. Given a sound ensembling approach, we would expect to see improvement as more prompts are added to the ensemble (up to some point). To study this question, we record results using $k\in\{1,5,10,20\}$ different rephrasings (randomly chosen for 1000 trials when $k<20$). Results for the Spatial, Goal, and Object task suites are shown in Figure 9. We see that the algorithm behaves favorably, where calibration error generally decreases as more instructions are included in the ensemble.
B.4 Calibration Across Task Time
Figures 10, 11, 12, 13, and 14 offer additional results on calibration across the task time horizon using the OpenVLA model. Figures 15, 16, and 17 offer results for the same experiment on UniVLA. Figure 18 shows further qualitative examples from the context-aware monitoring experiment.
Figure 10: Empirical study of calibration error across task time horizon for the Spatial task suite and the Quant-8 OpenVLA model.
Figure 11: Empirical study of calibration error across task time horizon for the Object task suite and the full precision OpenVLA model.
Figure 12: Empirical study of calibration error across task time horizon for the Object task suite and the Quant-8 OpenVLA model.
Figure 13: Empirical study of calibration error across task time horizon for the Goal task suite and the full precision OpenVLA model.
Figure 14: Empirical study of calibration error across task time horizon for the Goal task suite and the Quant-8 OpenVLA model.
Figure 15: Empirical study of calibration error across task time horizon for the Spatial task suite and the UniVLA model.
Figure 16: Empirical study of calibration error across task time horizon for the Object task suite and the UniVLA model.
Figure 17: Empirical study of calibration error across task time horizon for the Goal task suite and the UniVLA model.
B.4.1 Qualitative Examples of Context-Aware Confidence Monitoring
Additional qualitative examples of our confidence-aware monitoring demonstration are provided in Figure 18. These include cases where the strategy succeeds and others where it fails (e.g., confidence falls below the threshold while grasping in an ultimately successful trial).
Figure 18: Qualitative examples of a context-aware confidence calibration strategy for the task “put the wine bottle on the rack”. The red dashed line represents the 10% quantile of the confidence estimates output by the model across the task horizon, representing a potential threshold below which the robot may abstain from performing the task.
B.5 Calibration Across Action Dimensions
Figure 19 shows additional results for the action-wise recalibration experiments.
Figure 19: On the left of each pair of plots, we compare miscalibration across action dimensions. On the right, we compare the performance of typical Platt scaling, temperature scaling, histogram binning, and Platt binning to action-wise Platt scaling.